### v2

# Large-scale Long-tailed Disease Diagnosis on Radiology Images

Abstract
In this study, we aim to investigate the problem of large-scale, large-vocabulary disease classification for radiologic images, which can be formulated as a multi-modal, multi-anatomy, multi-label, long-tailed classification. Our main contributions are three folds: (i), on dataset construction, we build up an academically accessible, large-scale diagnostic dataset that encompasses 5568 disorders linked with 930 unique ICD-10-CM codes, containing 39,026 cases (192,675 scans). (ii), on model design, we present a novel architecture that enables to process arbitrary number of input scans, from various imaging modalities, which is trained with knowledge enhancement to leverage the rich domain knowledge; (iii), on evaluation, we initialize a new benchmark for multi-modal multi-anatomy long-tailed diagnosis. Our method shows superior results on it. Additionally, our final model serves as a pre-trained model, and can be finetuned to benefit diagnosis on various external datasets.
1 Introduction
In the ever-evolving landscape of clinical medicine, the advent of radiology techniques, such as X-ray, CT, MRI, and ultrasound, has truly revolutionized the medical field, offering a non-invasive yet deeply revealing perspective of the human body for disease diagnosis and management. These imaging techniques are now at the cusp of a new era with the integration of artificial intelligence (AI).
Recent literature highlights the significant potential of developing diagnostic models in the medical field, the developments can be generally cast into two categories: one is a specialist, which has already shown success in identifying and managing a wide array of diseases [15, 78, 86]. However, a notable limitation of these models is their specialization, as they often focus on a narrow range of disease categories, targeting limited anatomical regions, based on specific imaging modalities. This specialization restricts their ability to fully address the diverse and complex cases encountered in real-world clinical settings.
While on the other extreme, there is an emerging trend towards developing Generalist Medical Artificial Intelligence (GMAI) models [56, 81, 85]. Taking inspiration from the breakthroughs in natural language processing and computer vision, these models aim to amalgamate data from diverse sources, including various imaging methods, patient histories, and current medical research, to offer more comprehensive diagnostic and treatment solutions. Nonetheless, the development of GMAI models faces formidable challenges, such as the need for substantial computational power, meticulously curated multimodal datasets covering an extensive range of medical conditions and patient demographics, and advanced models equipped to tackle unique medical intricacies, like extremely unbalanced data distribution and the need for domain-specific expertise.
In this paper, we consider the problem of large-scale, large-vocabulary disease classification for radiologic images, marking a transition phase between specialist and generalist models. Specifically, compared to existing specialists, we aim to initiate the research for developing a computational model that can handle multi-modal, multi-anatomy, and multi-label disease diagnosis, in the face of extremely unbalanced distribution, embracing a wider scope of diagnoses across various anatomical regions and imaging modalities. In contrast to generalist models, our investigation offers a more feasible and targeted playground for exploring sophisticated algorithms in academic labs, offering opportunities for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs. Overall, we make contributions from three aspects, namely, a large-scale open dataset and its construction pipeline, preliminary model architecture exploration, and an evaluation benchmark.
On dataset construction, we build up an academically accessible, large-scale diagnostic dataset derived from Radiopaedia [5]. Each sample is associated to the International Classification of Diseases, i.e., ICD-10-CM [1], indicating the diagnostic category, for example, ‘S86’ refers to ‘Injury of muscle, fascia and tendon at lower leg level’. The dataset naturally displays an unbalanced distribution, varying from 1 to 964 cases in each disease category. Additionally, each case within the dataset involves multiple multi-modal scans, with the number of modalities ranging from 1 to 5, the number of images per case spanning between 1 and 30. As a result, we construct a long-tailed, multi-scan medical disease classification dataset, as shown in Figure 1, with 39,026 cases (192,675 scans) across 7 human anatomy regions and 9 diverse modalities covering 930 ICD-10-CM codes, 5568 disorders [^1] Disorders encompass a range of conditions such as abnormality, syndromes, injuries, poisonings, signs, symptoms, findings and diseases. Diseases are specific pathological conditions characterized by a set of identifiable signs and symptoms., termed as Radiopaedia3D Diagnosis Dataset (RP3D-DiagDS). We will release all the data, complete with corresponding disorders, ICD-10-CM codes, and detailed definitions.
On architecture design, we demonstrate a new model that supports both 2D and 3D input from various modalities, together with a transformer-based fusion module for comprehensive diagnosis. Specifically, for visual encoding, we explore two variants of backbones, namely, ResNet-based, and ViT-based to perform 2D or 3D unified encoding. Then we fuse the multi-scan information with a transformer-based fusion module, treating each image embedding as an input token. At training time, we adopt knowledge-enhanced training strategy [79, 84], specifically, we leverage the rich domain knowledge to pre-train a knowledge encoder with natural language and use it to guide the visual representation learning for disease diagnosis.
On evaluation, we carry out a series of ablation studies on the effectiveness of different training configurations, for example, visual backbones (ViT or ResNet), augmentation implementation, and depth of 3D input volumes. Then, we evaluate the model on our proposed benchmark of multi-modal multi-scan long-tailed multi-label diagnosis and demonstrate the superiority of our proposed methods on it. Furthermore, our trained model showed strong transferring abilities. By fine-tuning, it can benefit numerous external datasets, regardless of their image dimensions, imaging modalities, and shooting anatomies.
Figure 1: Overview of RP3D-DiagDS. There are 39,026 cases (192,675 scans) across 7 human anatomy regions and 9 diverse modalities covering 930 ICD-10-CM codes.
2 Related Work
2.1 Disease Classification Datasets
Open-source datasets play a crucial role in the development of AI for medical image analysis. Unlike natural images, the curation of medical image datasets presents unique challenges due to factors, such as privacy concerns, requirement of specialized domain knowledge for annotation, etc. In the literature, there has been a number of open-source datasets for disease diagnosis. Notably, large-scale datasets such as NIH ChestX-ray [76], MIMIC-CXR [38], and CheXpert [34] stand out for their comprehensive collection of annotated X-ray images, facilitating extensive research and advancements in automated disease classification. However, it’s important to note three key limitations. First, these open-source diagnosis datasets mainly consist of consist of chest X-rays [63, 12, 76, 34, 50, 66, 31, 36, 71, 25, 59, 22, 38, 19], thus are 2D images. There are only a few 3D datasets available, with a limited number of volumes [54, 3, 2, 4, 39, 43, 51, 33, 58, 66, 10, 7, 9]. Second, a significant portion of these large-scale datasets focuses solely on binary classification of specific diseases [71, 25, 31, 36]. Third, existing large-scale datasets containing a broad range of disease categories in different granularities [12, 82, 53], for example, in PadChest [12], disorders exist at different levels, including infiltrates, interstitial patterns, and reticular interstitial patterns. Such variation in granularity poses a great challenge for representation learning. In summary, these datasets fall short of meeting real-world clinical needs, which often involve complex, multimodal, and multi-image data from single patient. Therefore, constructing a dataset that mirrors the intricacies of actual clinical scenarios is necessary.
2.2 Specialist Diagnosis Models
The prevailing paradigm in earlier diagnostic models is a specialized model trained on a limited range of disease categories. These models focus on specific imaging modalities and are targeted towards particular anatomical regions. Specifically, ConvNets are widely used in medical image classification due to its outstanding performance, for example, [21, 69, 73, 29, 19, 46, 13, 74] have demonstrated excellent results on identifying a wide range of diseases. Recently, Vision Transformer (ViT) has garnered immense interest in the medical imaging community, numerous innovative approaches  [20, 80, 27, 70, 52, 87, 83, 65, 84, 67, 79, 48, 72, 26, 55] have emerged, leveraging ViTs as a foundation for further advancements in this field.
2.3 Generalist Medical Foundation Models
The other stream of work is generalist medical foundation models  [75, 81, 56, 85]. These models represent a paradigm shift in medical AI, aiming to create versatile, comprehensive AI systems capable of handling a wide range of tasks across different medical modalities, by leveraging large-scale, diverse medical data. MedPaLM M [75] reaches performance competitive with or exceeding the state-of-the-art (SOTA) on various medical benchmarks, demonstrating the potential of the generalist foundation model in disease diagnosis and beyond. RadFM [81] is the first medical foundation model capable of processing 3D multi-image inputs, demonstrating the versatility and adaptability of generalist models in processing complex imaging data. However, the development of GMAI models requires substantial computational power, which is often impractical for exploring sophisticated algorithms in academic labs.
2.4 Long-tailed Classification Strategy
Medical image diagnosis inherently faces long-tail challenges, as the prevalence of common diseases is usually substantially higher than that of rare diseases. A straightforward solution to the imbalance problem is to perform re-sampling during training [24, 28, 42, 49], e.g., over-sampling on tailed classes or under-sampling on head classes. However, this approach often triggers over-fitting in tail classes, resulting insufficient training for the head classes. Loss re-weighting is another widely used solution to tackle the long-tailed distribution problem [47, 68]. Focal loss [45, 47] refines cross-entropy by assigning lower weights to easily learnable data, thereby prioritizing challenging or misclassifiable instances. These approaches have primarily been explored on relatively small datasets with a limited number of categories. In this paper, our objective is to initiate the problem for large-scale, long-tailed, multi-scan medical disease classification.
3 Dataset Construction
In this section, we present the details of our dataset, RP3D-DiagDS, that follows a similar procedure as RP3D [81]. Specifically, cases in our dataset are sourced from the Radiopaedia website [5] – a growing peer-reviewed educational radiology resource website, that allows the clinicians to upload 3D volumes to better reflect real clinical scenarios. Additionally, all privacy issues have already been resolved by the clinicians at uploading time. It is worth noting that, unlike RP3D [81] that contains paired free-form text description and radiology scans for visual-language representation learning, here, we focus on multi-modal, multi-anatomy, and multi-label disease diagnosis (classification) under extremely unbalanced distribution.
Overall, the proposed dataset contains 39,026 cases, of 192,675 images from 9 diverse imaging modalities and 7 human anatomy regions, note that, each case may contain images of multiple scans. The data covers 5,568 different disorders, that have been manually mapped into 930 ICD-10-CM [1] codes.
3.1 Data Collection
Here, we describe the procedure of our data curation process, shown in Figure 2. Specifically, we collect three main components from each case on Radiopaedia webpage, namely, ‘Patient data’, ‘Radiology images’, and ‘Articles’  (example webpage shown in Figure 2 (a)). ‘Patient data’ includes brief information of this patient, for example, age, gender, etc. ‘Radiology images’ denote a series of radiology examination scans. Figure 5 (a) provides a statistical analysis on the number of images within one single case. ’Articles’ contains links to related articles named with corresponding disorders, which are treated as diagnosis labels and have been meticulously peer-reviewed by experts in Radiopaedia Editorial Board [^2] https://radiopaedia.org/editors.
Figure 2: The data collection pipeline. In (a) we show two original cases. Our dataset presents a multi-scan multi-modal multi-label long-tailed diagnosis task with 2D and 3D scan involed. In (b) we show the general data curation progress. We totally collect 39026 cases with 192675 images labeled by 5568 disorder classes and 930 ICD-10-CM classes. In addition to the intermediate case number, we also show the intermediate image number together with the final normal and abnormal case number in (b).
To start with, we can collect 50,970 cases linked with 10,670 articles from Radiopaedia website. However, naively adopting the article titles as disorders may lead to ambiguities from two aspects: (i) not all articles are related to disorders, (ii) article titles can be written with different granularity, for example, “pneumonia” and “bacterial pneumonia” can be referred to as different disorders, though they should ideally be arranged in a hierarchical structure, (iii) normal cases from Radiopaedia are not balanced in modalities and anatomies. Next, we discuss 3-stage procedure to alleviate the above-mentioned challenges.
Article Filtering. We leverage the GPT-4 [64] to automatically filter the article list, leaving those refer to disorders. Specifically, taking inspiration from the self-consistency prompts [77], we design two different query prompts with similar meanings, as shown in Figure 3 (a). An article name is labeled as disorder if GPT-4 gives consistent positive results from both prompts, while for those GPT-4 gives inconsistent results, we do manually check them. To measure the quality for filtering, we randomly sample a portion of data for manual checking to control its quality. The confusion metrics are shown in Figure 3 (c). The $100\%$ precision score indicates our auto-filtering strategy, can strictly ensure the left ones to be disorders. Eventually, 5342 articles can pass the first auto-criterion and 226 pass the second manually checking, resulting in 5,568 disorder classes, as shown in Figure 2 (b). Cases not linked to any articles are excluded, ultimately yielding 38,858 cases.
Figure 3: The article filtering pipeline. In (a), we show the two GPT prompts and (b) is the whole pipeline where we leverage the two prompts to query GPT-4 twice and only the consistent results will be used for auto-filtering, otherwise, we will check it manually. (c) shows the accuracy of the auto-filtering cases. Note that, all the 200 article samples evaluated in (c) have a consistent prediction under the two GPT-4 prompts.
Mapping disorders to ICD-10-CM. Upon getting disorder classes, they may fall into varying hierarchical granularity levels, here, we hope to map them to internationally recognized standards. For this purpose, we utilize the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes, which is a customized version of the ICD-10 used for coding diagnoses in the U.S. healthcare system. The ICD was originally designed as a health care classification system, providing a system of diagnostic codes for classifying diseases, including nuanced classifications of a wide variety of signs, symptoms, abnormal findings, complaints, social circumstances, and external causes of injury or disease. Specifically, we hire ten medical PhD students to manually map the article titles into ICD-10-CM codes. After cross-checking by a ten-year clinician, we have mapped the 5,568 disorders into the corresponding ICD-10-CM codes. We unify various disorders into the first hierarchical level of ICD-10-CM code tree, resulting in 930 ICD-10-CM classes. For example, both “J12.0 Adenoviral pneumonia” and “J12 Viral pneumonia” are mapped to the code “J12 Viral pneumonia, not elsewhere classified”, dismissing the ambiguities caused by diagnosis granularity. Note that, we will provide the ICD-10-CM codes along with the original 5,568 classes, each accompanied by its corresponding definition, thus, the dataset can be employed for training diagnosis or visual-language models.
Adding Normal Cases. Despite there are normal cases on Radiopaedia [^3] https://radiopaedia.org/articles/normal-imaging-examples, covering most of anatomies and modalities, we additionally collect more normal cases from MISTR [^4] https://mistr.usask.ca/odin/, with images available for research under a CC-BY-NC-SA 4.0 license. The expanded normal cases include 168 cases. The distribution of modalities and anatomical regions will be shown in the data statistics section later.
Finally, we get 39,026 cases containing 192,675 images labeled by 5,568 disorder classes and 930 ICD-10-CM classes and will continually maintain the dataset, growing the case number.
3.2 Dataset Statistics
In this section, we provide detailed analysis on our proposed dataset, from three aspects, namely, modality coverage, anatomy coverage, and disease coverage.
Table 1: Comparison between RP3D-DiagDS and existing disease diagnosis datasets. #Image refers to the number of images in the dataset. #Disorders represents the number of categories of labeled disorders. 2D/3D indicates the dimension of images in the dataset, and the case-level column indicates whether the dataset contains multiple images per patient. Our dataset demonstrates significant increase in both size and diversity.
Dataset Anatomy Modality #Image #Disorders 2D 3D Case-level
Brain-Tumor [3] Head and Neck MRI 7k 4
CE-MRI  [17] Head and Neck MRI 3k 3
Brain-Tumor-17  [4] Head and Neck MRI 4.4k 17
OASIS[39, 43, 51] Head and Neck MRI 3.2k 2
ICH2020 [33] Head and Neck CT 82 5
COVID19-CT-DB [58] Chest CT 1k 9
Covid-CXR2 [66] Chest X-ray 19k 3
POCUS [10] Chest Ultrasound 1.1k 3
MIA-COV19 [40] Chest CT 5.3k 2
LIDC-IDRI [7] Chest CT 1.0k 3
NIH Chestxray [76] Chest X-ray 100k 14
PadChest [12] Chest X-ray 160k 193
CheXpert [34] Chest X-ray 224k 14
RSNA [71] Chest X-ray 30k 2
SIIM-ACR [25] Chest X-ray 15k 2
MIMIC-CXR [38] Chest X-ray 371k 14
VinDr-CXR [59] Chest X-ray 18k 28
VinDr-PCXR [63] Chest X-ray 9k 51
VinDr-Mammo [61] Breast Mammography 20k 10
DDSM [32, 44] Breast Mammography 55k 5
CMMD2022 [14] Breast Mammography 5.2k 2
BUSI [6] Breast Ultrasound 780 3
VinDr-SpineXr [62] Spine X-ray 10k 13
MRNet [9] Knee MRI 1370 3
DeepLesion [82] Whole body Radiology 32k 22
RadImageNet [53] Whole body CT, MRI, Ultrasound 1.35M 165
RP3D-DiagDS (ours) Whole body Radiology 192k 5568 (ICD-10-CM:930)*
    We map disorders to ICD-10-CM for standardization.
Analysis on Modality Coverage. RP3D-DiagDS comprises images from 9 modalities, namely, computed tomography (CT), magnetic resonance imaging (MRI), X-ray, Ultrasound, Fluoroscopy, Nuclear medicine, Mammography, DSA (angiography), and Barium Enema. Each case may include images from multiple modalities, to ensure precise and comprehensive diagnosis of disorders. Overall, approximately 19.4% of the cases comprise images from two modalities, while around 2.9% involve images from three to five modalities. The remaining cases are associated with image scans from a single modality. The distribution of modalities among all abnormal samples is illustrated in Figure 4 (a). The modalities of normal cases follow similar distributions with the abnormal cases, as shown in Figure 4 (b).
(a) Abnormal Cases Distribution in Modality
(b) Normal Cases Distribution in Modality
(c) Abnormal Cases Distribution in Anatomy
(d) Normal Cases Distribution in Anatomy
Figure 4: The distribution of imaging modalities of abnormal (a) and normal (b) cases in RP3D-DiagDS and anatomies of abnormal (c) and normal (d) cases in RP3D-DiagDS. Each label is annotated with the class name, number of cases, and the corresponding proportion.
Analysis on Anatomy Coverage. RP3D-DiagDS comprises images from various anatomical regions, including the head and neck, spine, chest, breast, abdomen and pelvis, upper limb, and lower limb, providing comprehensive coverage of the entire human body. The statistics are shown in Figure 4 (c) and 4 (d).
Figure 5: Distribution Analysis. In (a), we show the distribution for the number of images in one case. Each case may include multiple images from patient history scans, different modalities, different angles or conditions. In (b), we demonstrate the long-tailed distributions for disorder and ICD-10-CM classes. We also categorize these classes into three categories: “head class”, “body class” and “tail class” based on the number of cases. Notably, to better show the main part of the case distributions in (b), we clip the axes, indicated by the dotted axes lines.
Analysis on Disease Coverage. For both disorder and disease classification, each case can correspond to multiple disorders, resulting in RP3D-DiagDS a multi-label classification dataset. As shown in Figure 5 (b), the distributions exhibit extremely long-tailed pattern, rendering such 2D & 3D image classification problem as a long-tailed multi-label classification task. We define the ‘head class’ category with case counts greater than 100, the ‘body class’ category with case counts between 30 and 100, and the ‘tail class’ category with case counts less than 30.
4 Method
In this section, we aim to initiate a preliminary investigation on computational architectures for large-scale, long-tailed disease diagnosis on radiology, specifically, we start by defining the problem of case-level multi-label classification in Sec. 4.1, then we elaborate the architecture in Sec. 4.2, and knowledge-enhanced training strategy in Sec. 4.3.
4.1 Problem Formulation
We consider one case with its disease labels, represented as $\mathcal{X}=\{x_{1},\dots,x_{S},y_{1},\dots,y_{c}\}$, where $x_{i}$ denotes a specific scan (2D or 3D) under a certain radiologic examination, for example, MRI, CT or X-ray within $\mathcal{X}$, $S$ denotes the scan number for the case and $c$ represents the total diagnostic labels. Note that, there may be multiple scans from the same or different modalities for each patient.
To illustrate this concept, consider a case where a patient undergoes a physical examination, including a chest CT and a chest X-ray as shown in Figure 6. In this instance, $S=2$ reflects the two scans involved in this case. If the patient is diagnosed with nodule and pneumothorax, the labels for these diseases are set to 1, reflecting their presence in the patient’s diagnosis.
Our goal is to train a model that can solve the above multi-class, multi-scan diagnosis problem:
$\mathcal{Y}=\Phi(\mathcal{X})=\Phi_{\text{cls}}(\Phi_{\text{fuse}}(\Phi_{\text{visual}}(x_{1}),\Phi_{\text{visual}}(x_{2}),\cdots,\Phi_{\text{visual}}(x_{S})))\in\mathcal{R}^{c},$ (1)
the model is composed of a visual encoder $\Phi_{\text{visual}}$, a fusion module $\Phi_{\text{fuse}}$ and a classifier $\Phi_{\text{cls}}$.
4.2 Architecture
Figure 6: The architecture of our proposed visual encoder and fusion module, together with the knowledge enhancement strategy. (a) shows the details of the vision encoder. We design two variants to fit in the two main visual backbones, i.e., ResNet and ViT. (b) shows the transformer-based fusion module, enabling case-level information fusion. (c) shows the knowledge enhancement strategy. We first pre-train a text encoder with extra medical knowledge with contrastive learning, i.e., synonyms, descriptions and hierarchy, termed as knowledge encoder and then we view the text embedding as a natural classifier to guide the diagnosis classification.
Our proposed architecture consists of two key components: (i) a visual encoder that processes 2D or 3D input scan; (ii) a transformer-based fusion module, merging all information to perform case-level diagnosis.
4.2.1 Visual Encoder
We consider two popular variants of the visual encoder, namely, ResNet [30] and ViT [23], as shown in Figure 6 (a). The visual encoding progress can generally be formulated as:
$v_{i}=\Phi_{\text{visual}}(x_{i})\in\mathcal{R}^{d}.$ (2)
where $x_{i}\in\mathcal{R}^{C\times H\times W\times(D)}$ denotes the input scan, $C,H,W$ refer to the image channel, height and width of the input scan respectively. $D$ is optional, and only available for 3D input scans. The main challenge for architecture design comes from the requirement to process scans in both 2D and 3D formats. Here, we train separate normalisation modules to convert the 2D or 3D inputs into feature maps of same resolution, and further passed into shared encoding module.
ResNet. For 3D scans, they are first fed into 3D ResNet, followed by average pooling on depth to aggregate the information of the extra dimension; while for 2D scans, they are fed to the 2D ResNet to perform the same down-sampling ratio on height and width as for 3D. After normalization, both 2D and 3D scans are transformed into feature maps with same resolution, $f_{i}\in\mathcal{R}^{d_{\text{res}}\times h\times w}$, where $d_{\text{res}}$ is the intermediate feature dimension and $h,w$ denote the normalized size. Then, the feature maps are passed into a shared ResNet to get the final visual embedding.
ViT. For 3D scans, we convert the input volume into a series of non-overlapped 3D cubes, and pass them into MLP projection layers, to get vector embeddings; while for 2D scans, the input scan is broken into 2D patches, and projected into vector embeddings with another set of MLP layers. As ViT enables to handle sequences of variable tokens, we can now pass the resulting vectors into a shared ViT to get the final visual embedding. For position embedding, we adopt two sets of learnable position embeddings, one for 2D input, the other for 3D, further indicating the network about the input dimenstion.
4.2.2 Fusion Module
For case-level diagnosis, we propose to aggregate information from multiple scans with a trainable module. As shown in Figure 6 (b), we adopt transformer encoders, specifically, we first initialize a set of learnable modality embeddings, denoted as $\{p_{1},\dots,p_{M}\}$, where $M$ denotes the total number of possible imaging modalities. Given certain visual embedding ($v_{i}$) from modality $j$, we first add the corresponding modality embedding ($p_{j}$) to it, indicating which radiologic modality it is from. Then we feed all visual embeddings from one case into the fusion module, and output the fused visual embedding $v_{\text{fuse}}\in\mathcal{R}^{d}$, from the “[cls]” token, similar to paper [23], denoted as:
$v_{\text{fuse}}=\Phi_{\text{fuse}}(\Phi_{\text{visual}}(x_{1}),\Phi_{\text{visual}}(x_{2}),\cdots,\Phi_{\text{visual}}(x_{S}))\in\mathcal{R}^{d}.$ (3)
Till here, we have computed the case-level visual embedding, which can be passed into classifier for disease diagnosis. Here, we adopt a knowledge-enhanced training strategy [79, 84], that has shown to be superior in long-tailed recognition problem, as detailed in the following section.
4.3 Knowledge-enhanced Training
With knowledge enhancement, we hope to leverage the rich domain knowledge in medicine to enhance the long-tailed classification. Our key insight is that the long-tailed diseases may fundamentally have some shared symptoms or radiologic pathologies with the head classes, which could be encoded in text format.
In detail, we first pre-train a text encoder with medical knowledge, termed as knowledge encoder, where names of similar disorder or diseases are projected to similar embeddings, for example, ‘lung disease’ is closer to ‘pneumonia’ in the embedding space than ‘brain disease’. Then we freeze the text embeddings, and use them to guide the training of vision encoder, as shown in Figure 6 (c).
Knowledge Encoder Pre-training. We leverage several extra knowledge bases to pre-train a knowledge encoder, including Radiopaedia, ICD10-CM, and UMLS. Specifically, for each disorder term, we collect its definitions, radiologic features from Radiopaedia articles, synonyms, clinical information, and hierarchy structure from ICD10-CM, as well as definitions from UMLS. We aim to train the text encoder with the following considerations:
  • Synonyms. If two terms are identified as similar synonyms, we expect them to be close in text embedding space, like ‘Salmonella enteritis’ and ‘Salmonella gastroenteritis’.
  • Hierarchy. In the context of hierarchical relationships, we expect that if a disease is a fine-grained classification of another disease, its embeddings should be closer than those of unrelated diseases. For example, ‘J93, Pneumothorax and air leak’ should exhibit closer embeddings with ‘J93.0, Spontaneous tension pneumothorax’ than with ‘J96.0, Acute respiratory failure’.
  • Descriptions. For terms associated with descriptions, radiologic features or clinical information, we expect their embeddings to be close in the text embedding space. For example, ‘Intestinal Neuroendocrine Tumor’ and ‘A well-differentiated, low or intermediate grade tumor with neuroendocrine differentiation that arises from the small or large intestine’.
We start from an off-the-shelf text encoder, namely, MedCPT-Query-Encoder [37], and adopt contrastive learning for further finetuning. Given a target medical terminology name encoded by the text encoder, denoted as $f_{\text{tar}}$, its corresponding medical texts, i.e., synonyms, containing terminologies or related descriptions are treated as positive cases. Similarly, we encode them with the text encoder, denoted as $f^{+}$, and other non-related text embedding is treated as negative cases $f^{-}$. Note that, we keep positive and negative cases consistent in format, e.g., when the positive case is related to description sentence, the negative cases are always some non-related description sentences instead of the short name words. The final objective can be formulated as:
$\mathcal{L}_{\text{knowledge}}=-\log\frac{e^{f_{\text{tar}}^{T}\cdot f^{+}/\tau}}{\sum_{n=1}^{N}e^{f_{\text{tar}}^{T}\cdot f^{-}_{n}/\tau}},$ (4)
where $\tau$ refers to the temperature, and $N$ denotes total sampled negative cases. By optimizing the contrastive loss we can further finetune the text encoder, resulting in a knowledge encoder, termed as $\Phi_{k}$
Knowledge-guided Classification. After training the knowledge encoder, we use it to encode the disorder/disease names into text embeddings, for example, denoting the names as $\{T_{1},T_{2},\dots,T_{c}\}$, where $T_{j}$ is a disorder/disease name like “pneumonia” or “lung tumor”. We embed these free texts with the knowledge encoder as:
$t_{j}=\Phi_{k}(T_{j})\in\mathcal{R}^{d},$ (5)
where $t_{j}$ denotes the text embedding. The resulting text embeddings are used for case-level diagnosis:
$p=v_{\text{fuse}}\cdot t\in\mathcal{R}^{c},$ (6)
where $p$ is the final result. We use classical binary cross entropy (BCE) loss as the final training objective, denoted as:
$\mathcal{L}=-\sum_{i=1}^{c}\mathcal{Y}_{i}\cdot\log(p_{i})+(1-\mathcal{Y}_{i})\cdot\log(1-p_{i}).$ (7)
5 Experiments
In this section, we will introduce our evaluation settings. Specifically, we first establish a benchmark for case-level multi-modal, multi-scan, long-tailed disorder/disease diagnosis. Second, we treat RP3D-DiagDS as a large-scale dataset for pre-training, and evaluate its transferring ability to various existing datasets.
5.1 Long-tailed Classification on RP3D-DiagDS
With the proposed dataset, we consider the problem as a multi-label classification task under a long-tailed distribution. In this section, we provide an overview of the training and evaluation protocols employed in our study. We first introduce the dataset split, followed by details of evaluation metrics for assessment.
5.1.1 Dataset Split
We randomly split our dataset into training (train and validation) and test sets in a (7:1):2 ratio. We treat the class with [100:][100:][ 100 : ] positive cases as head classes, $[30,100)$ cases as medium classes, and [:30)[:30)[ : 30 ) cases as tail classes. Consequently, from a total of 38,858 cases, the training set comprises 27,201 cases, and the test set includes 7,772 cases. Following on the dataset construction procedure, we perform classification task on two class sets at different granularities: (i) 5568 disorders + normal; (ii) 930 ICD-10-CM codes + normal. On different sets, we will have different head/medium/tail class set following our definition. Note that, the normal class is always treated as one of the head classes. Next, we will only talk about the abnormal classes.
Disorders. At the disorder level, there are 85 head classes, 470 medium classes, and 5014 tail classes. Notably, as the case number of some tailed classes is extremely small, some classes may only appear in training or test sets, but not both. This issue only happens to the tail classes. It is important to note that our knowledge encoder enables to get embedding for unseen diseases, ensuring that evaluation is not affected by such challenge. As a result, there are 5,305/3,399 classes in our training and testing split, respectively.
ICD-10-CM. At ICD-10-CM level, disorders with similar meanings have been mapped to same codes, for example, ‘tuberculosis’, ‘primary pulmonary tuberculosis’ all correspond to ‘A15 respiratory tuberculosis’. This merge operation results in more cases in one class, consequently more head classes. Specifically, we have 165 head classes, 230 medium classes, and 536 tail classes. Similarly, some tail classes may only be in training or testing split, resulting in 902/759 classes for our final training and testing splits, respectively.
It is important to point that while we ensure the inclusion of head and medium classes in both the training and test sets, some tail classes are inevitably exclusive from certain divisions, which is treated as our future work to collect more samples of such disorders. We report the results separately for head/medium/tail classes, while focusing primarily on the head classes.
5.1.2 Evaluation Metrics
In this section, we describe the evaluation metrics in detail. Note that, the following metrics can all be calculated per class. For multi-class cases, we all use macro-average on classes to report the scores by default. For example, “AUC” for multi-class classification denotes the “Macro-averaged AUC on classes”.
AUC. Area Under Curve [11] denotes the area under ROC (receiver operating characteristic) curve. This has been widely used in medical diagnosis, due to its clinical meanings and robustness in unbalanced distribution.
AP. Average Precision (AP) is calculated as the weighted mean of precisions at each threshold. Specifically, for each class, we rank all samples according to the prediction score, then shift the threshold to a series of precision-recall points and draw the precision-recall (PR) curve, AP equals the area under the curve. This score measures whether the unhealthy samples are ranked higher than healthy ones. We report Mean Average Precision (mAP), which is the average of AP of each class.
F1 and MCC. F1 score is the harmonic mean of the precision and recall. It is widely used in diagnosis tasks. MCC [18] denotes the Matthews Correlation Coefficient metric. It ranges from $-1$ to $1$ and can be calculated as follow:
$\text{MCC}=\frac{\text{TN}\times\text{TP}-\text{FN}\times\text{FP}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}$
Both metrics need a specific threshold to compute, and we choose one by maximizing F1 on the validation set following former papers [59, 79, 80, 84].
Recall@FPR. We also report the recall scores at different false positive rate (FPR), i.e., sample points from class-wise ROC curves. Specifically, We report the Recall@0.01, Recall@0.05, Recall@0.1 score, denoting the recall scores at 0.01, 0.05, 0.1 FPR respectively.
5.2 Transfer Learning to External Datasets
In addition to evaluating on our own benchmark, we also consider transferring our final model to other external datasets, demonstrating its transferring abilities on image distribution shift and label space change. In the following sections, we start by introducing the used external datasets, that cover various medical imaging modalities and anatomies. Then we detail the fine-tuning settings.
5.2.1 External Datasets
We choose external evaluation dataset with the following principles:
  • Imaging Modalities. We hope to cover most radiologic modalities in our external evaluation, e.g., 2D X-ray, 2D CT/MRI slices and 3D CT/MRI scans, demonstrating our model can benefit all of them;
  • Human Anatomies. We hope to cover many human anatomies, e.g., brain, head and neck, chest, spine abdomen and limb, demonstrating our model can benefit all of them;
  • Label Space. We hope to cover two cases, i.e., seen and unseen classes. Seen classes refer to those labels that have appeared in RP3D-DiagDS. Conversely, unseen extra classes denote those not included.
As a result, we pick the following datasets, and report AUC scores to compare with others on these external evaluation.
  • CXR14 [76] is a widely-used chest X-ray (2D) diagnosis dataset containing 112,120 frontal-view X-ray images of 30,805 (collected from the year 1992 to 2015) unique patients with 14 finding labels. We follow its official split and evaluate the SOTA [80] on the split.
  • VinDr-Spine [60] is a spine X-ray (2D) diagnosis dataset comprising 10,469 images from 5,000 studies. We follow K-Diag [79] to use the 8 unique finding labels and the official split.
  • VinDr-Mammo [60] is a mammography (2D) diagnosis dataset comprising 20,000 images (5,000 four-view scans). Each scan was manually annotated with a 5-level BI-RADS score. We view this as a multi-class classification task with the official split.
  • ADNI [35] is a 3D brain MRI dataset focused on Alzheimer’s disease comprising 112141 images, including AD, MCI, CN and other classifications. We random split it follow 63,846/15,962 for training and testing and reproduce the state-of-the-art method on it.
  • MosMedData [57] is a 3D chest CT dataset on 5-level COVID-19 grading comprising 1,110 images. We follow the official split. We random split it follow 888/222 for training and testing and reproduce the state-of-the-art method on it.
5.2.2 Fine-tuning Diagnosis
Our final model can serve as a pre-trained model and be fine-tuned on each downstream task to improve the final performance, demonstrating the merit of our dataset. Specifically, for the dataset with single image input, we simply adopt the pre-trained visual encoder module, i.e., discarding the fusion module. While for the datset with multi image input, we will use them all. In both cases, the final classification layer will be trained from scratch. In addition to using all available external training data, we also consider to use 1%, 10%, 30% portion data for few-shot learning.
5.3 Implementation Details
At training time, we consider two diagnose tasks in different granularities: Disorder-level classification (5569 classes) and ICD10-level classification (931 Classes). The image input will all be resized to $512\times 512\times D$ in height, width and depth respectively. For 3D scans, the depth is treated as a factor for ablation study, and will be discussed in experiment section. In vision encoder, we adopt two separated modules for normalization and a shared module to compute the final embedding. The detailed architecture design will also be further discussed in our ablation study. In fusion module, we use a 6-layer transformer encoder with learnable ‘[cls]’ token for final prediction.
For augmentation, we adopt Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation, implemented from the MONAI [16] package. For optimization, we utilize the AdamW optimizer with a cosine learning rate curve, setting the maximum learning rate to $lr=1\times e^{-5}$, with an adjustable batch size ranging from 4 to 32 depending on the input image depth and model scale to avoid out-of-memory error. The total training duration spans 100 epochs, with the initial 5 epochs for warm-up. At fine-tuning stage, we adopt the similar model architecture and optimization setting.
6 Results
We conduct an ablation study to identify the optimal architecture and hyper-parameters for our model. Then, we present the evaluation results, focusing on the ICD-10-CM classification and disorders classification. Lastly, we use our dataset for large-scale pre-training, and finetune it on various external datasets, to demonstrate the model’s transferring ability.
6.1 Ablation Study
To explore the optimal model architecture and parameter configurations, including the fusion strategy, visual encoder architecture, the depth of 3D scans, and augmentation, we conduct a series of ablation studies on a subset of original dataset, comprising 200 disorder categories with most cases, termed as SubSet@200. In the default experiment setting, we use 16 as the 3D scan depth, and ResNet as visual backbone, without knowledge enhancement and augmentation strategy. While conducting ablation study on certain component, we keep other setting unchanged.
6.1.1 Visual Encoder Architecture
In this section, we investigate different backbone architectures for diagnosis, specifically, we compare the ResNet-based and the ViT-based models. As shown in Table 2, “Seperated” denotes the separated visual encoder to perform 2D and 3D normalization and “Shared” denotes the shared encoder part for both 2D and 3D scans. We make two observations, (i) ResNet-based architecture performs better than ViT-based ones, (ii) increasing the capacity of shared encoder is more beneficial, e.g., ResNet-34+ResNet-18 vs. ResNet-18+ResNet-34. (iii) Deeper ResNet structure has very limited improvement, e.g., ResNet-34+ResNet-34 vs. ResNet-18+ResNet-34.
Table 2: Ablation study on various settings of visual encoder architectures. “Normalisation” denotes the separated visual encoder part to perform 2D and 3D normalization and “Shared Enc.” denotes the shared encoder part for both 2D and 3D scans.
Visual Encoder Architecture AUC AP F1 MCC R@F0.01 R@F0.05 R@D0.1
Normalisation Shared Enc.
VIT 2-layer MLP 6-layer VIT 79.98 6.01 13.57 14.69 17.78 33.56 47.21
4-layer MLP 6-layer VIT 80.57 6.13 13.49 14.78 17.68 34.01 47.44
2-layer MLP 12-layer VIT 81.69 6.40 14.71 15.30 18.11 34.73 48.84
4-layer MLP 12-layer VIT 82.03 6.67 14.94 15.66 18.20 34.99 49.52
ResNet ResNet-18 ResNet-18 86.91 11.00 16.77 18.63 20.42 41.87 59.38
ResNet-34 ResNet-18 86.99 11.15 17.14 19.21 20.82 44.67 61.13
ResNet-18 ResNet-34 87.06 11.27 17.36 19.23 21.48 44.38 61.54
ResNet-34 ResNet-34 87.10 11.31 17.66 19.41 21.33 44.19 62.25
6.1.2 Image Dimension
Here, we aim to investigate the effect of input resolution, i.e., increasing the depth of input volumes. Due to the constraints of GPU physical memory, we experiment with 16, 24, and 32 as the depths for 3D input volume. We employ trilinear interpolation to resample 3D scans to the same size.
Table 3 illustrates the performance by varying the depths of 3D images. It can be observed that an increase in depth can bring clear performance gain, suggesting that the detailed depth information is critical to perform diagnosis. Consequently, with more slices, the model tend to yield more favorable results.
Table 3: Ablation study on different 3D scan depth size settings. Considering the computational cost, we evaluated the commonly seen 16, 24, 32 depth sizes on our datsets.
3D Image Depth AUC AP F1 MCC R@0.01 R@0.05 R@0.1
16 87.06 11.27 17.36 19.23 21.48 44.38 61.54
24 87.49 11.41 17.84 19.39 21.76 45.72 62.53
32 88.13 11.98 18.02 19.67 22.84 47.46 63.91
6.1.3 Augmentation
Here, we evaluate the effectiveness of augmentation on our dataset. Specifically, we adopt four augmentation strategies with $15\%$ probability each, namely, Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation. As a result, about half of the training data will be applied at least one augmentation in each training batch. As shown in Table 4, adopting data augmentation shows notable performance improvement. This improvement is particularly evident on AUC and AP when employing the ViT as the visual backbone while, still, ResNet-based model performs better.
Table 4: Ablation on adopting augmentation strategy. In the table, we show the effectiveness of adopting reasonable data augmentation for our task. “ResNet” denotes the ResNet-based visual encoder architecture and “ViT” denotes the ViT-basde one.
Augmentation Visual Encoder AUC AP F1 MCC R@F0.01 R@F0.05 R@F0.1
YES ResNet 88.96 13.92 22.04 23.08 25.75 51.61 66.47
ViT 84.22 8.69 16.36 17.18 19.83 38.70 54.04
NO ResNet 87.06 11.27 17.36 19.23 21.48 44.38 61.54
ViT 81.69 6.40 14.71 15.30 18.11 34.73 48.84
6.1.4 Summary
In conclusion, we find the ResNet-based model with augmentation strategy and unifying the 3D scan depth to 32 are the most suitable setting for our long-tailed case-level multi-modal diagnosis task, which will be used for training on the entire dataset in the following sections.
6.2 Evaluation on Rad3D-DiagDS
In this section, we train our model on the entire Rad3D-DiagDS training set, and evaluate on the test split respectively, as shown in Table 5, and the AUC curve comparison is shown in Figure 7.
Specifically, we carry out the experiments at two levels, i.e., disorder and ICD-10-CM classes. We denote fusion module as FM for short and knowledge enhancement as KE. In cases where without FM or KE, we adopt max pooling on the predictions of different images from the same case (Check supplementary for more details), serving as a baseline. Then, we add the fusion module and knowledge enhancement step-by-step, to improve the model’s performance.
Table 5: Classification results on Disorders and ICD-10-CM levels. In the table “FM” represents the fusion module and “KE” represents the knowledge enhancement strategy. We report the results on Head/Medium/Tail class sets separately.
Granularity Class Methods Metrics
FM
KE
AUC AP F1 MCC R@0.01 R@0.05 R@0.1
Disorders Head
91.91 13.56 22.08 23.66 30.88 60.50 74.70
93.56 14.2 24.77 25.84 36.19 65.21 79.83
94.24 15.33 25.68 26.71 37.06 66.55 81.37
Medium
93.62 9.19 17.16 20.42 28.24 59.32 73.80
94.03 11.47 19.18 23.14 30.04 63.89 78.69
94.69 12.38 20.64 24.07 31.52 65.34 78.73
Tail
88.55 4.34 8.03 12.96 9.13 27.46 41.43
90.01 8.97 9.35 13.30 10.04 27.02 42.32
90.64 9.25 9.89 14.38 10.98 27.98 43.53
ICD-10-CM Head
88.29 12.01 19.80 20.85 22.95 50.12 62.25
89.55 13.89 21.92 23.41 25.04 52.19 67.26
90.89 14.37 22.67 24.29 26.11 53.82 69.16
Medium
88.37 7.85 15.58 18.02 22.88 49.62 65.25
89.90 8.98 17.33 19.89 24.37 51.33 64.93
91.67 9.56 18.32 20.77 25.68 52.85 66.63
Tail
85.28 4.02 7.53 11.71 7.94 21.28 34.41
86.34 4.77 8.81 12.85 8.83 22.36 37.11
86.55 4.81 8.69 12.74 8.86 22.75 37.97
(a) ROCs on Disorders (Head).
(b) ROCs on Disorders (Medium).
(c) ROCs on Disorders (Tail).
(d) ROCs on ICD-10-CM (Head).
(e) ROCs on ICD-10-CM (Medium).
(f) ROCs on ICD-10-CM (Tail).
Figure 7: ROC curves on Disorders and ICD-10-CM, including head/medium/tail parts respectively. The shadow in the figure shown the 95% CI (Confidence interval) and FM, KE are short for Fusion Module and Knowledge Enhancement.
Discussion on Fusion Module. To demonstrate the effectiveness of fusion module, we start with a baseline without FM or KE. As shown in Table 5, adding the fusion module can greatly improve the results on Head, Medium and Tail classes at both disorder and ICD-10-CM levels, showing the critical role for case-level information fusion in diagnosis task. These results align well with our expectations. In clinical practice, the examination of one modality for a diagnosis is often insufficient. A thorough and meticulous diagnostic process typically involves an integrated review of all test results. Each test is weighted differently, depending on how its results correspond with other tests. Our fusion model adeptly mirrors this comprehensive approach, demonstrating its effectiveness in simulating the nuanced process of clinical diagnosis.
Discussion on Knowledge Enhancement. As shown in Table 5, incorporating knowledge enhancement further promotes the final diagnosis performance on both disorder and ICD-10-CM cases, validating the assumption that a better text embedding space trained by rich domain knowledge can greatly help the visual representation learning.
Discussion on ROC Curves. The solid lines in the Figure 7 represents the median AUC value. This value is derived from a process where, for each category, 1000 random samples are taken and the median AUC value is calculated. This procedure is repeated multiple times (1000), with the final curve representing the median of these values. The ROC curve is illustrated by thess solid lines. Accompanying each solid lines is shaded area, which denotes the 95% confidence interval (CI). For each type of implementation, this shaded area is significantly higher than the portion above the solid line. This implies that the AUC value represented by the solid line is generally higher than the class-wise average across different data splits. As shown in the figure, the highest AUC value has reached 0.96, and the lowest value is 0.86, this observed pattern suggests, there exists a few categories that are more challenging to learn than others.
[…]
    The numbers are borrowed from the referred papers. We are strictly align with them in train and test split.
6.3 External Evaluation
In this section, we explore the transferring ability of our final model, where we finetune the model on different downstream datasets across various imaging modalities, anatomies and target classes. As shown in Table 6.2, we can see significant performance improvement on various datasets with different data portions, compring to models trained from scratch. Additionally, in most cases, our model can also surpass former SOTAs significantly without any task-specific designs, e.g., architectures, loss functions, demonstrating that our RP3D-DiagDS can also serve as a superior large-scale supervised pre-training dataset for medical domain.
7 Clinical Impact
In this paper, we consider the problem of multi-modality, multi-anatomy, long-tailed disease diagnosis, which has great practical clinical meanings.
First, in clinical practice, patients may get multiple radiologic examinations during their long-term treatment progress from various medical departments. However, existing work on disease classification accepting one image scan, can hardly handle such circumstance, hindering the development of accurate and comprehensive diagnostic models. Our work compensate this and better meet the practical scenarios.
Second, comparing with existing diagnosis works that focus on common general disease classes, our efforts on long-tailed rare classes are more vital for clinical usage. For common diseases, AI diagnosis system, usually, can only help accelerate diagnosis procedure for clinicians, while the hint on rare classes is also critical.
Third, our final model can support finetuning on external diagnosis task. In clinical application, local medical centers can leverage this superiority even with a few training samples. This is remarkable especially considering that, for rare diseases, only a very few cases could be accessible in reality.
8 Limitation
Despite the effectiveness of our proposed dataset, and architecture, there remains some improvements: First, on model design, in fusion step, we can use more image tokens to represent a scan rather than a pooled single vector. The latter may lead to excessive loss of image information during fusion. The model size can be further increased to investigate the effect of model capacities; Second, new loss functions should be explored to tackle such large-scale, long-tailed disorder/disease classification task. Third, on disorder to ICD-10-CM mapping progress, the annotators are labeled on class-name level, i.e., only disorder names are provided, causing some ambiguous classes unable to find strict corresponding ICD-10-CM codes. Though we mark out this class in our shared data files, if providing more case-level information, the mapping could be more accurate. We treat these as future work.
9 Conclusion
In this paper, we focus on solving the problem of multi-modal, multi-label long-tailed case-level diagnosis task. Specifically, we propose a new large-scale diagnosis dataset, namely, RP3D-DiagDS, with 39,026 cases (192,675 scans) labeled with detailed disorders covering 930 ICD-10-CM. On model design, we propose one unified architecture that supports both 2D and 3D input, together with a fusion module, to integrate information from multiple scans of one patient. Additionally, we adopt knowledge enhancement training strategy, leveraging the rich medical domain knowledge to improve the radiologic diagnosis performance. Our final train model also shows strong transferring ability to various external datasets regardless of their imaging modalities, shooting anatomies and target classes. We believe this work can serve as a transition phase between the specialist and generalist models, offering a more feasible and targeted playground for exploring sophisticated algorithms in academia labs, providing a unique opportunity for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs.
10 Data and Code Availability
Our dataset RP3D-DiagDS can be found in https://huggingface.co/datasets/QiaoyuZheng/RP3D-DiagDS and our codes can be found in https://github.com/qiaoyu-zheng/RP3D-Diag

### v3

@@
 Abstract
-In this study, we aim to investigate the problem of large-scale, large-vocabulary disease classification for radiologic images, which can be formulated as a multi-modal, multi-anatomy, multi-label, long-tailed classification.
-Our main contributions are three folds: (i), on dataset construction, we build up an academically accessible, large-scale diagnostic dataset that encompasses 5568 disorders linked with 930 unique ICD-10-CM codes, containing 39,026 cases (192,675 scans).
-(ii), on model design, we present a novel architecture that enables to process arbitrary number of input scans, from various imaging modalities, which is trained with knowledge enhancement to leverage the rich domain knowledge; (iii), on evaluation, we initialize a new benchmark for multi-modal multi-anatomy long-tailed diagnosis.
-Our method shows superior results on it.
-Additionally, our final model serves as a pre-trained model, and can be finetuned to benefit diagnosis on various external datasets.
+Developing a generalist radiology diagnosis system can greatly enhance clinical diagnostics.
+In this paper, we introduce RadDiag, a foundational model supporting 2D and 3D inputs across various modalities and anatomies, using a transformer-based fusion module for comprehensive disease diagnosis.
+Due to patient privacy concerns and the lack of large-scale radiology diagnosis datasets, we utilize high-quality, clinician-reviewed radiological images available online with diagnosis labels.
+Our dataset, RP3D-DiagDS, contains 40,936 cases with 195,010 scans covering 5,568 disorders (930 unique ICD-10-CM codes).
+Experimentally, our RadDiag achieves 95.14% AUC on internal evaluation with the knowledge-enhancement strategy.
+Additionally, RadDiag can be zero-shot applied or fine-tuned to external diagnosis datasets sourced from various hospitals, demonstrating state-of-the-art results.
+In conclusion, we show that publicly shared medical data on the Internet is a tremendous and valuable resource that can potentially support building a generalist AI for healthcare.
 1 INTRODUCTION
-In the ever-evolving landscape of clinical medicine, the advent of radiology techniques, such as X-ray, CT, MRI, and ultrasound, has truly revolutionized the medical field, offering a non-invasive yet deeply revealing perspective of the human body for disease diagnosis and management.
-These imaging techniques are now at the cusp of a new era with the integration of artificial intelligence (AI).
-Recent literature highlights the significant potential of developing diagnostic models in the medical field, the developments can be generally cast into two categories: one is a specialist, which has already shown success in identifying and managing a wide array of diseases [15, 78, 86].
-However, a notable limitation of these models is their specialization, as they often focus on a narrow range of disease categories, targeting limited anatomical regions, based on specific imaging modalities.
-This specialization restricts their ability to fully address the diverse and complex cases encountered in real-world clinical settings.
-While on the other extreme, there is an emerging trend towards developing Generalist Medical Artificial Intelligence (GMAI) models [56, 81, 85].
-Taking inspiration from the breakthroughs in natural language processing and computer vision, these models aim to amalgamate data from diverse sources, including various imaging methods, patient histories, and current medical research, to offer more comprehensive diagnostic and treatment solutions.
-Nonetheless, the development of GMAI models faces formidable challenges, such as the need for substantial computational power, meticulously curated multimodal datasets covering an extensive range of medical conditions and patient demographics, and advanced models equipped to tackle unique medical intricacies, like extremely unbalanced data distribution and the need for domain-specific expertise.
-In this paper, we consider the problem of large-scale, large-vocabulary disease classification for radiologic images, marking a transition phase between specialist and generalist models.
-Specifically, compared to existing specialists, we aim to initiate the research for developing a computational model that can handle multi-modal, multi-anatomy, and multi-label disease diagnosis, in the face of extremely unbalanced distribution, embracing a wider scope of diagnoses across various anatomical regions and imaging modalities.
-In contrast to generalist models, our investigation offers a more feasible and targeted playground for exploring sophisticated algorithms in academic labs, offering opportunities for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs.
-Overall, we make contributions from three aspects, namely, a large-scale open dataset and its construction pipeline, preliminary model architecture exploration, and an evaluation benchmark.
-On dataset construction, we build up an academically accessible, large-scale diagnostic dataset derived from Radiopaedia [5].
-Each sample is associated to the International Classification of Diseases, i.e., ICD-10-CM [1], indicating the diagnostic category, for example, ‘S86’ refers to ‘Injury of muscle, fascia and tendon at lower leg level’.
-The dataset naturally displays an unbalanced distribution, varying from 1 to 964 cases in each disease category.
-Additionally, each case within the dataset involves multiple multi-modal scans, with the number of modalities ranging from 1 to 5, the number of images per case spanning between 1 and 30.
-As a result, we construct a long-tailed, multi-scan medical disease classification dataset, as shown in Figure 1, with 39,026 cases (192,675 scans) across 7 human anatomy regions and 9 diverse modalities covering 930 ICD-10-CM codes, 5568 disorders [^1] Disorders encompass a range of conditions such as abnormality, syndromes, injuries, poisonings, signs, symptoms, findings and diseases.
-Diseases are specific pathological conditions characterized by a set of identifiable signs and symptoms., termed as Radiopaedia3D Diagnosis Dataset (RP3D-DiagDS).
-We will release all the data, complete with corresponding disorders, ICD-10-CM codes, and detailed definitions.
-On architecture design, we demonstrate a new model that supports both 2D and 3D input from various modalities, together with a transformer-based fusion module for comprehensive diagnosis.
-Specifically, for visual encoding, we explore two variants of backbones, namely, ResNet-based, and ViT-based to perform 2D or 3D unified encoding.
-Then we fuse the multi-scan information with a transformer-based fusion module, treating each image embedding as an input token.
-At training time, we adopt knowledge-enhanced training strategy [79, 84], specifically, we leverage the rich domain knowledge to pre-train a knowledge encoder with natural language and use it to guide the visual representation learning for disease diagnosis.
-On evaluation, we carry out a series of ablation studies on the effectiveness of different training configurations, for example, visual backbones (ViT or ResNet), augmentation implementation, and depth of 3D input volumes.
-Then, we evaluate the model on our proposed benchmark of multi-modal multi-scan long-tailed multi-label diagnosis and demonstrate the superiority of our proposed methods on it.
-Furthermore, our trained model showed strong transferring abilities.
-By fine-tuning, it can benefit numerous external datasets, regardless of their image dimensions, imaging modalities, and shooting anatomies.
+In the landscape of clinical medicine, the emergence of radiology techniques, for example, X-ray, CT, and MRI, has truly revolutionized the medical field, which enables to diagnose by reading radiological images.
+Recent literature has demonstrated significant potential for developing AI models for disease diagnosis in the medical field [83, 72].
+However, a notable limitation of the existing models is their specialization, as these ‘specialist’ models often focus on a narrow range of disease categories, and thus usually fail to accommodate the breadth and complexity of clinical scenarios.
+Conversely, an ideal radiology diagnosis system is expected to comprehensively analyze the combination of arbitrary historical scans across any anatomical regions and imaging modalities for a wide range of diseases.
+In this paper, we aim to build a foundational radiology diagnostic model, that supports information fusion from an arbitrary number of 2D and 3D inputs across different modalities and anatomic sites.
+Generally speaking, there are three major challenges.
+First, lack of datasets for training.
+Despite the large volume of radiology data generated per day across the world, publicly available ones are extremely limited due to privacy issues.
+Existing open-source datasets are normally collected from a certain modality or anatomy with very limited diseases, thus significantly hindering the development of powerful disease diagnosis models.
+Second, lack of unified architecture formulation.
+Models for diagnosing radiological diseases are often tailored to particular datasets, which are optimized for a narrow spectrum of diseases linked with particular imaging modalities [22, 34].
+In contrast, developing a foundation model for disease diagnosis requires the architecture to be capable of fusing information from an arbitrary number of images of various modalities, for example, X-ray, CT, and MRI.
+Third, lack of benchmark to monitor progress.
+Benchmarking the models’ performance on disease diagnosis relies on datasets that well simulate real-world scenarios, for instance, containing many diseases, long-tail distribution, etc.
+Existing benchmarks typically focus on a very limited number of diseases, with single modality inputs [11, 51].
+To comprehensively measure the progress in disease diagnosis, a high-quality benchmark has yet to be established.
 Figure 1: Overview of RP3D-DiagDS. There are 39,026 cases (192,675 scans) across 7 human anatomy regions and 9 diverse modalities covering 930 ICD-10-CM codes.
-2 Related Work
-2.1 Disease Classification Datasets
-Open-source datasets play a crucial role in the development of AI for medical image analysis.
-Unlike natural images, the curation of medical image datasets presents unique challenges due to factors, such as privacy concerns, requirement of specialized domain knowledge for annotation, etc.
-In the literature, there has been a number of open-source datasets for disease diagnosis.
-Notably, large-scale datasets such as NIH ChestX-ray [76], MIMIC-CXR [38], and CheXpert [34] stand out for their comprehensive collection of annotated X-ray images, facilitating extensive research and advancements in automated disease classification.
-However, it’s important to note three key limitations.
-First, these open-source diagnosis datasets mainly consist of consist of chest X-rays [63, 12, 76, 34, 50, 66, 31, 36, 71, 25, 59, 22, 38, 19], thus are 2D images.
-There are only a few 3D datasets available, with a limited number of volumes [54, 3, 2, 4, 39, 43, 51, 33, 58, 66, 10, 7, 9].
-Second, a significant portion of these large-scale datasets focuses solely on binary classification of specific diseases [71, 25, 31, 36].
-Third, existing large-scale datasets containing a broad range of disease categories in different granularities [12, 82, 53], for example, in PadChest [12], disorders exist at different levels, including infiltrates, interstitial patterns, and reticular interstitial patterns.
-Such variation in granularity poses a great challenge for representation learning.
-In summary, these datasets fall short of meeting real-world clinical needs, which often involve complex, multimodal, and multi-image data from single patient.
-Therefore, constructing a dataset that mirrors the intricacies of actual clinical scenarios is necessary.
-2.2 Specialist Diagnosis Models
-The prevailing paradigm in earlier diagnostic models is a specialized model trained on a limited range of disease categories.
-These models focus on specific imaging modalities and are targeted towards particular anatomical regions.
-Specifically, ConvNets are widely used in medical image classification due to its outstanding performance, for example, [21, 69, 73, 29, 19, 46, 13, 74] have demonstrated excellent results on identifying a wide range of diseases.
-Recently, Vision Transformer (ViT) has garnered immense interest in the medical imaging community, numerous innovative approaches  [20, 80, 27, 70, 52, 87, 83, 65, 84, 67, 79, 48, 72, 26, 55] have emerged, leveraging ViTs as a foundation for further advancements in this field.
-2.3 Generalist Medical Foundation Models
-The other stream of work is generalist medical foundation models  [75, 81, 56, 85].
-These models represent a paradigm shift in medical AI, aiming to create versatile, comprehensive AI systems capable of handling a wide range of tasks across different medical modalities, by leveraging large-scale, diverse medical data.
-MedPaLM M [75] reaches performance competitive with or exceeding the state-of-the-art (SOTA) on various medical benchmarks, demonstrating the potential of the generalist foundation model in disease diagnosis and beyond.
-RadFM [81] is the first medical foundation model capable of processing 3D multi-image inputs, demonstrating the versatility and adaptability of generalist models in processing complex imaging data.
-However, the development of GMAI models requires substantial computational power, which is often impractical for exploring sophisticated algorithms in academic labs.
-2.4 Long-tailed Classification Strategy
-Medical image diagnosis inherently faces long-tail challenges, as the prevalence of common diseases is usually substantially higher than that of rare diseases.
-A straightforward solution to the imbalance problem is to perform re-sampling during training [24, 28, 42, 49], e.g., over-sampling on tailed classes or under-sampling on head classes.
-However, this approach often triggers over-fitting in tail classes, resulting insufficient training for the head classes.
-Loss re-weighting is another widely used solution to tackle the long-tailed distribution problem [47, 68].
-Focal loss [45, 47] refines cross-entropy by assigning lower weights to easily learnable data, thereby prioritizing challenging or misclassifiable instances.
-These approaches have primarily been explored on relatively small datasets with a limited number of categories.
-In this paper, our objective is to initiate the problem for large-scale, long-tailed, multi-scan medical disease classification.
-3 Dataset Construction
-In this section, we present the details of our dataset, RP3D-DiagDS, that follows a similar procedure as RP3D [81].
-Specifically, cases in our dataset are sourced from the Radiopaedia website [5] – a growing peer-reviewed educational radiology resource website, that allows the clinicians to upload 3D volumes to better reflect real clinical scenarios.
-Additionally, all privacy issues have already been resolved by the clinicians at uploading time.
-It is worth noting that, unlike RP3D [81] that contains paired free-form text description and radiology scans for visual-language representation learning, here, we focus on multi-modal, multi-anatomy, and multi-label disease diagnosis (classification) under extremely unbalanced distribution.
-Overall, the proposed dataset contains 39,026 cases, of 192,675 images from 9 diverse imaging modalities and 7 human anatomy regions, note that, each case may contain images of multiple scans.
-The data covers 5,568 different disorders, that have been manually mapped into 930 ICD-10-CM [1] codes.
-3.1 Data Collection
-Here, we describe the procedure of our data curation process, shown in Figure 2.
-Specifically, we collect three main components from each case on Radiopaedia webpage, namely, ‘Patient data’, ‘Radiology images’, and ‘Articles’  (example webpage shown in Figure 2 (a)).
-‘Patient data’ includes brief information of this patient, for example, age, gender, etc.
-‘Radiology images’ denote a series of radiology examination scans.
-Figure 5 (a) provides a statistical analysis on the number of images within one single case.
-’Articles’ contains links to related articles named with corresponding disorders, which are treated as diagnosis labels and have been meticulously peer-reviewed by experts in Radiopaedia Editorial Board [^2] https://radiopaedia.org/editors.
-Figure 2: The data collection pipeline.
-In (a) we show two original cases.
-Our dataset presents a multi-scan multi-modal multi-label long-tailed diagnosis task with 2D and 3D scan involed.
-In (b) we show the general data curation progress.
-We totally collect 39026 cases with 192675 images labeled by 5568 disorder classes and 930 ICD-10-CM classes.
-In addition to the intermediate case number, we also show the intermediate image number together with the final normal and abnormal case number in (b).
-To start with, we can collect 50,970 cases linked with 10,670 articles from Radiopaedia website.
-However, naively adopting the article titles as disorders may lead to ambiguities from two aspects: (i) not all articles are related to disorders, (ii) article titles can be written with different granularity, for example, “pneumonia” and “bacterial pneumonia” can be referred to as different disorders, though they should ideally be arranged in a hierarchical structure, (iii) normal cases from Radiopaedia are not balanced in modalities and anatomies.
-Next, we discuss 3-stage procedure to alleviate the above-mentioned challenges.
-Article Filtering.
-We leverage the GPT-4 [64] to automatically filter the article list, leaving those refer to disorders.
-Specifically, taking inspiration from the self-consistency prompts [77], we design two different query prompts with similar meanings, as shown in Figure 3 (a).
-An article name is labeled as disorder if GPT-4 gives consistent positive results from both prompts, while for those GPT-4 gives inconsistent results, we do manually check them.
-To measure the quality for filtering, we randomly sample a portion of data for manual checking to control its quality.
-The confusion metrics are shown in Figure 3 (c).
-The $100\%$ precision score indicates our auto-filtering strategy, can strictly ensure the left ones to be disorders.
-Eventually, 5342 articles can pass the first auto-criterion and 226 pass the second manually checking, resulting in 5,568 disorder classes, as shown in Figure 2 (b).
-Cases not linked to any articles are excluded, ultimately yielding 38,858 cases.
-Figure 3: The article filtering pipeline.
-In (a), we show the two GPT prompts and (b) is the whole pipeline where we leverage the two prompts to query GPT-4 twice and only the consistent results will be used for auto-filtering, otherwise, we will check it manually.
-(c) shows the accuracy of the auto-filtering cases.
-Note that, all the 200 article samples evaluated in (c) have a consistent prediction under the two GPT-4 prompts.
-Mapping disorders to ICD-10-CM.
-Upon getting disorder classes, they may fall into varying hierarchical granularity levels, here, we hope to map them to internationally recognized standards.
-For this purpose, we utilize the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes, which is a customized version of the ICD-10 used for coding diagnoses in the U.S.
-healthcare system.
-The ICD was originally designed as a health care classification system, providing a system of diagnostic codes for classifying diseases, including nuanced classifications of a wide variety of signs, symptoms, abnormal findings, complaints, social circumstances, and external causes of injury or disease.
-Specifically, we hire ten medical PhD students to manually map the article titles into ICD-10-CM codes.
-After cross-checking by a ten-year clinician, we have mapped the 5,568 disorders into the corresponding ICD-10-CM codes.
-We unify various disorders into the first hierarchical level of ICD-10-CM code tree, resulting in 930 ICD-10-CM classes.
-For example, both “J12.0 Adenoviral pneumonia” and “J12 Viral pneumonia” are mapped to the code “J12 Viral pneumonia, not elsewhere classified”, dismissing the ambiguities caused by diagnosis granularity.
-Note that, we will provide the ICD-10-CM codes along with the original 5,568 classes, each accompanied by its corresponding definition, thus, the dataset can be employed for training diagnosis or visual-language models.
-Adding Normal Cases.
-Despite there are normal cases on Radiopaedia [^3] https://radiopaedia.org/articles/normal-imaging-examples, covering most of anatomies and modalities, we additionally collect more normal cases from MISTR [^4] https://mistr.usask.ca/odin/, with images available for research under a CC-BY-NC-SA 4.0 license.
-The expanded normal cases include 168 cases.
-The distribution of modalities and anatomical regions will be shown in the data statistics section later.
-Finally, we get 39,026 cases containing 192,675 images labeled by 5,568 disorder classes and 930 ICD-10-CM classes and will continually maintain the dataset, growing the case number.
-3.2 Dataset Statistics
-In this section, we provide detailed analysis on our proposed dataset, from three aspects, namely, modality coverage, anatomy coverage, and disease coverage.
-Table 1: Comparison between RP3D-DiagDS and existing disease diagnosis datasets.
-#Image refers to the number of images in the dataset.
-#Disorders represents the number of categories of labeled disorders.
-2D/3D indicates the dimension of images in the dataset, and the case-level column indicates whether the dataset contains multiple images per patient.
-Our dataset demonstrates significant increase in both size and diversity.
-Dataset Anatomy Modality #Image #Disorders 2D 3D Case-level
-Brain-Tumor [3] Head and Neck MRI 7k 4
-CE-MRI  [17] Head and Neck MRI 3k 3
-Brain-Tumor-17  [4] Head and Neck MRI 4.4k 17
-OASIS[39, 43, 51] Head and Neck MRI 3.2k 2
-ICH2020 [33] Head and Neck CT 82 5
-COVID19-CT-DB [58] Chest CT 1k 9
-Covid-CXR2 [66] Chest X-ray 19k 3
-POCUS [10] Chest Ultrasound 1.1k 3
-MIA-COV19 [40] Chest CT 5.3k 2
-LIDC-IDRI [7] Chest CT 1.0k 3
-NIH Chestxray [76] Chest X-ray 100k 14
-PadChest [12] Chest X-ray 160k 193
-CheXpert [34] Chest X-ray 224k 14
-RSNA [71] Chest X-ray 30k 2
-SIIM-ACR [25] Chest X-ray 15k 2
-MIMIC-CXR [38] Chest X-ray 371k 14
-VinDr-CXR [59] Chest X-ray 18k 28
-VinDr-PCXR [63] Chest X-ray 9k 51
-VinDr-Mammo [61] Breast Mammography 20k 10
-DDSM [32, 44] Breast Mammography 55k 5
-CMMD2022 [14] Breast Mammography 5.2k 2
-BUSI [6] Breast Ultrasound 780 3
-VinDr-SpineXr [62] Spine X-ray 10k 13
-MRNet [9] Knee MRI 1370 3
-DeepLesion [82] Whole body Radiology 32k 22
-RadImageNet [53] Whole body CT, MRI, Ultrasound 1.35M 165
-RP3D-DiagDS (ours) Whole body Radiology 192k 5568 (ICD-10-CM:930)*
-    We map disorders to ICD-10-CM for standardization.
-Analysis on Modality Coverage.
-RP3D-DiagDS comprises images from 9 modalities, namely, computed tomography (CT), magnetic resonance imaging (MRI), X-ray, Ultrasound, Fluoroscopy, Nuclear medicine, Mammography, DSA (angiography), and Barium Enema.
-Each case may include images from multiple modalities, to ensure precise and comprehensive diagnosis of disorders.
-Overall, approximately 19.4% of the cases comprise images from two modalities, while around 2.9% involve images from three to five modalities.
-The remaining cases are associated with image scans from a single modality.
-The distribution of modalities among all abnormal samples is illustrated in Figure 4 (a).
-The modalities of normal cases follow similar distributions with the abnormal cases, as shown in Figure 4 (b).
-(a) Abnormal Cases Distribution in Modality
-(b) Normal Cases Distribution in Modality
-(c) Abnormal Cases Distribution in Anatomy
-(d) Normal Cases Distribution in Anatomy
-Figure 4: The distribution of imaging modalities of abnormal (a) and normal (b) cases in RP3D-DiagDS and anatomies of abnormal (c) and normal (d) cases in RP3D-DiagDS.
-Each label is annotated with the class name, number of cases, and the corresponding proportion.
-Analysis on Anatomy Coverage.
-RP3D-DiagDS comprises images from various anatomical regions, including the head and neck, spine, chest, breast, abdomen and pelvis, upper limb, and lower limb, providing comprehensive coverage of the entire human body.
-The statistics are shown in Figure 4 (c) and 4 (d).
-Figure 5: Distribution Analysis.
-In (a), we show the distribution for the number of images in one case.
-Each case may include multiple images from patient history scans, different modalities, different angles or conditions.
-In (b), we demonstrate the long-tailed distributions for disorder and ICD-10-CM classes.
-We also categorize these classes into three categories: “head class”, “body class” and “tail class” based on the number of cases.
-Notably, to better show the main part of the case distributions in (b), we clip the axes, indicated by the dotted axes lines.
-Analysis on Disease Coverage.
-For both disorder and disease classification, each case can correspond to multiple disorders, resulting in RP3D-DiagDS a multi-label classification dataset.
-As shown in Figure 5 (b), the distributions exhibit extremely long-tailed pattern, rendering such 2D & 3D image classification problem as a long-tailed multi-label classification task.
-We define the ‘head class’ category with case counts greater than 100, the ‘body class’ category with case counts between 30 and 100, and the ‘tail class’ category with case counts less than 30.
-4 Method
-In this section, we aim to initiate a preliminary investigation on computational architectures for large-scale, long-tailed disease diagnosis on radiology, specifically, we start by defining the problem of case-level multi-label classification in Sec.
-4.1, then we elaborate the architecture in Sec.
-4.2, and knowledge-enhanced training strategy in Sec.
-4.3.
-4.1 Problem Formulation
-We consider one case with its disease labels, represented as $\mathcal{X}=\{x_{1},\dots,x_{S},y_{1},\dots,y_{c}\}$, where $x_{i}$ denotes a specific scan (2D or 3D) under a certain radiologic examination, for example, MRI, CT or X-ray within $\mathcal{X}$, $S$ denotes the scan number for the case and $c$ represents the total diagnostic labels.
-Note that, there may be multiple scans from the same or different modalities for each patient.
+To overcome the above-mentioned challenges, we consider to exploit the widely available data on Internet, especially those focusing on high-quality radiology samples from professional institutions and educational websites, that have undergone meticulous de-identification, and reviewed by a panel of experienced clinicians.
+Specifically, we develop a flexible data collection pipeline, with each sample associated with an International Classification of Diseases code, i.e., ICD-10-CM [1], indicating the diagnostic category.
+The dataset presents a long-tailed distribution across disease categories.
+The case numbers range from 1 to 964 with some rare diseases.
+It includes multi-modal scans per case, with modalities varying from 1 to 5 and image counts between 1 and 30.
+As a result, we collect a multi-scan medical disease classification dataset, with 40,936 cases (195,010 scans) across 7 anatomy regions and 9 modalities, covering 930 ICD-10-CM codes and 5568 disorders, including a broad spectrum of medical conditions.
+Architecturally, we conduct preliminary explorations on the design that supports both 2D and 3D input from various modalities, together with a transformer-based fusion module for comprehensive diagnosis.
+Specifically, for visual encoding, we explore three variants of backbones, namely, ResNet-based, ViT-based and ResNet-ViT-mixing to perform 2D or 3D unified encoding.
+Then, we fuse the multi-scan information with a transformer-based fusion module, treating each image embedding as an input token.
+At training time, we adopt knowledge-enhanced training strategy [77, 83], where we first leverage the rich domain knowledge to pre-train a knowledge encoder with natural language, and use it to guide the visual representation learning for disease diagnosis.
+In experiments, we evaluate our model on a benchmark, constructed by images from online resources, featuring large-scale, multi-modal, multi-scan, long-tailed disease diagnosis.
+Additionally, we demonstrate that our proposed model can function as a general-purpose image encoder to capture a better image representation for radiology, making it possible to train and classify new sets via data-efficient fine-tuning or zero-shot transferring, leading to improved performance across diverse disease diagnosis tasks on numerous datasets, regardless of their image dimensions, modalities and anatomies.
+This paper demonstrates that publicly shared medical information is a tremendous resource that can be harnessed to advance the development of general AI for healthcare.
+2 RESULTS
+In this section, we start by briefly introducing our considered problem formulation, followed with showing the characteristics of our dataset for case-level multi-modal, multi-scan, long-tailed disorder/disease diagnosis.
+Then, we leverage a test split of the dataset to carry out a series of studies on diagnostic performance and establish a case-level diagnosis benchmark.
+Lastly, we train and assess the developed model accordingly, along with its transferring ability to various external datasets.
+Problem Formulation
+In this paper, our objective is to develop a radiology diagnostic model capable of integrating information from various images within a single case.
+We define a case as encompassing a collection of both 2D and 3D radiological images from different modalities and anatomical sites of a patient, who may be diagnosed with multiple conditions.
+More specifically, we represent a case and its associated disease labels as $\mathcal{X}=\{x_{1},\dots,x_{S},y_{1},\dots,y_{c}\}$, where each $x_{i}$ signifies a specific scan (either 2D or 3D) from a radiological examination such as MRI, CT, or X-ray.
+Here, $S$ is the number of scans for the case, and $c$ is the total number of diagnostic labels.
+It is noteworthy that a patient’s case might include multiple scans across the same or different modalities.
 To illustrate this concept, consider a case where a patient undergoes a physical examination, including a chest CT and a chest X-ray as shown in Figure 7.
@@
 $\mathcal{Y}=\Phi(\mathcal{X})=\Phi_{\text{cls}}(\Phi_{\text{fuse}}(\Phi_{\text{visual}}(x_{1}),\Phi_{\text{visual}}(x_{2}),\cdots,\Phi_{\text{visual}}(x_{S})))\in\mathcal{R}^{c},$ (1)
-the model is composed of a visual encoder $\Phi_{\text{visual}}$, a fusion module $\Phi_{\text{fuse}}$ and a classifier $\Phi_{\text{cls}}$.
-4.2 Architecture
-Figure 6: The architecture of our proposed visual encoder and fusion module, together with the knowledge enhancement strategy.
-(a) shows the details of the vision encoder.
-We design two variants to fit in the two main visual backbones, i.e., ResNet and ViT.
-(b) shows the transformer-based fusion module, enabling case-level information fusion.
-(c) shows the knowledge enhancement strategy.
-We first pre-train a text encoder with extra medical knowledge with contrastive learning, i.e., synonyms, descriptions and hierarchy, termed as knowledge encoder and then we view the text embedding as a natural classifier to guide the diagnosis classification.
-Our proposed architecture consists of two key components: (i) a visual encoder that processes 2D or 3D input scan; (ii) a transformer-based fusion module, merging all information to perform case-level diagnosis.
-4.2.1 Visual Encoder
-We consider two popular variants of the visual encoder, namely, ResNet [30] and ViT [23], as shown in Figure 6 (a). The visual encoding progress can generally be formulated as:
-$v_{i}=\Phi_{\text{visual}}(x_{i})\in\mathcal{R}^{d}.$ (2)
-where $x_{i}\in\mathcal{R}^{C\times H\times W\times(D)}$ denotes the input scan, $C,H,W$ refer to the image channel, height and width of the input scan respectively.
-$D$ is optional, and only available for 3D input scans.
-The main challenge for architecture design comes from the requirement to process scans in both 2D and 3D formats.
-Here, we train separate normalisation modules to convert the 2D or 3D inputs into feature maps of same resolution, and further passed into shared encoding module.
-ResNet.
-For 3D scans, they are first fed into 3D ResNet, followed by average pooling on depth to aggregate the information of the extra dimension; while for 2D scans, they are fed to the 2D ResNet to perform the same down-sampling ratio on height and width as for 3D.
-After normalization, both 2D and 3D scans are transformed into feature maps with same resolution, $f_{i}\in\mathcal{R}^{d_{\text{res}}\times h\times w}$, where $d_{\text{res}}$ is the intermediate feature dimension and $h,w$ denote the normalized size.
-Then, the feature maps are passed into a shared ResNet to get the final visual embedding.
-ViT.
-For 3D scans, we convert the input volume into a series of non-overlapped 3D cubes, and pass them into MLP projection layers, to get vector embeddings; while for 2D scans, the input scan is broken into 2D patches, and projected into vector embeddings with another set of MLP layers.
-As ViT enables to handle sequences of variable tokens, we can now pass the resulting vectors into a shared ViT to get the final visual embedding.
-For position embedding, we adopt two sets of learnable position embeddings, one for 2D input, the other for 3D, further indicating the network about the input dimenstion.
-4.2.2 Fusion Module
-For case-level diagnosis, we propose to aggregate information from multiple scans with a trainable module.
-As shown in Figure 6 (b), we adopt transformer encoders, specifically, we first initialize a set of learnable modality embeddings, denoted as $\{p_{1},\dots,p_{M}\}$, where $M$ denotes the total number of possible imaging modalities.
-Given certain visual embedding ($v_{i}$) from modality $j$, we first add the corresponding modality embedding ($p_{j}$) to it, indicating which radiologic modality it is from.
-Then we feed all visual embeddings from one case into the fusion module, and output the fused visual embedding $v_{\text{fuse}}\in\mathcal{R}^{d}$, from the “[cls]” token, similar to paper [23], denoted as:
-$v_{\text{fuse}}=\Phi_{\text{fuse}}(\Phi_{\text{visual}}(x_{1}),\Phi_{\text{visual}}(x_{2}),\cdots,\Phi_{\text{visual}}(x_{S}))\in\mathcal{R}^{d}.$ (3)
-Till here, we have computed the case-level visual embedding, which can be passed into classifier for disease diagnosis.
-Here, we adopt a knowledge-enhanced training strategy [79, 84], that has shown to be superior in long-tailed recognition problem, as detailed in the following section.
-4.3 Knowledge-enhanced Training
-With knowledge enhancement, we hope to leverage the rich domain knowledge in medicine to enhance the long-tailed classification.
-Our key insight is that the long-tailed diseases may fundamentally have some shared symptoms or radiologic pathologies with the head classes, which could be encoded in text format.
-In detail, we first pre-train a text encoder with medical knowledge, termed as knowledge encoder, where names of similar disorder or diseases are projected to similar embeddings, for example, ‘lung disease’ is closer to ‘pneumonia’ in the embedding space than ‘brain disease’.
-Then we freeze the text embeddings, and use them to guide the training of vision encoder, as shown in Figure 6 (c).
-Knowledge Encoder Pre-training.
-We leverage several extra knowledge bases to pre-train a knowledge encoder, including Radiopaedia, ICD10-CM, and UMLS.
-Specifically, for each disorder term, we collect its definitions, radiologic features from Radiopaedia articles, synonyms, clinical information, and hierarchy structure from ICD10-CM, as well as definitions from UMLS.
-We aim to train the text encoder with the following considerations:
-  • Synonyms. If two terms are identified as similar synonyms, we expect them to be close in text embedding space, like ‘Salmonella enteritis’ and ‘Salmonella gastroenteritis’.
-  • Hierarchy.
-In the context of hierarchical relationships, we expect that if a disease is a fine-grained classification of another disease, its embeddings should be closer than those of unrelated diseases.
-For example, ‘J93, Pneumothorax and air leak’ should exhibit closer embeddings with ‘J93.0, Spontaneous tension pneumothorax’ than with ‘J96.0, Acute respiratory failure’.
-  • Descriptions.
-For terms associated with descriptions, radiologic features or clinical information, we expect their embeddings to be close in the text embedding space.
-For example, ‘Intestinal Neuroendocrine Tumor’ and ‘A well-differentiated, low or intermediate grade tumor with neuroendocrine differentiation that arises from the small or large intestine’.
-We start from an off-the-shelf text encoder, namely, MedCPT-Query-Encoder [37], and adopt contrastive learning for further finetuning.
-Given a target medical terminology name encoded by the text encoder, denoted as $f_{\text{tar}}$, its corresponding medical texts, i.e., synonyms, containing terminologies or related descriptions are treated as positive cases.
-Similarly, we encode them with the text encoder, denoted as $f^{+}$, and other non-related text embedding is treated as negative cases $f^{-}$.
-Note that, we keep positive and negative cases consistent in format, e.g., when the positive case is related to description sentence, the negative cases are always some non-related description sentences instead of the short name words.
-The final objective can be formulated as:
-$\mathcal{L}_{\text{knowledge}}=-\log\frac{e^{f_{\text{tar}}^{T}\cdot f^{+}/\tau}}{\sum_{n=1}^{N}e^{f_{\text{tar}}^{T}\cdot f^{-}_{n}/\tau}},$ (4)
-where $\tau$ refers to the temperature, and $N$ denotes total sampled negative cases.
-By optimizing the contrastive loss we can further finetune the text encoder, resulting in a knowledge encoder, termed as $\Phi_{k}$
-Knowledge-guided Classification.
-After training the knowledge encoder, we use it to encode the disorder/disease names into text embeddings, for example, denoting the names as $\{T_{1},T_{2},\dots,T_{c}\}$, where $T_{j}$ is a disorder/disease name like “pneumonia” or “lung tumor”.
-We embed these free texts with the knowledge encoder as:
-$t_{j}=\Phi_{k}(T_{j})\in\mathcal{R}^{d},$ (5)
-where $t_{j}$ denotes the text embedding. The resulting text embeddings are used for case-level diagnosis:
-$p=v_{\text{fuse}}\cdot t\in\mathcal{R}^{c},$ (6)
-where $p$ is the final result. We use classical binary cross entropy (BCE) loss as the final training objective, denoted as:
-$\mathcal{L}=-\sum_{i=1}^{c}\mathcal{Y}_{i}\cdot\log(p_{i})+(1-\mathcal{Y}_{i})\cdot\log(1-p_{i}).$ (7)
-5 Experiments
-In this section, we will introduce our evaluation settings.
-Specifically, we first establish a benchmark for case-level multi-modal, multi-scan, long-tailed disorder/disease diagnosis.
-Second, we treat RP3D-DiagDS as a large-scale dataset for pre-training, and evaluate its transferring ability to various existing datasets.
-5.1 Long-tailed Classification on RP3D-DiagDS
-With the proposed dataset, we consider the problem as a multi-label classification task under a long-tailed distribution.
-In this section, we provide an overview of the training and evaluation protocols employed in our study.
-We first introduce the dataset split, followed by details of evaluation metrics for assessment.
-5.1.1 Dataset Split
-We randomly split our dataset into training (train and validation) and test sets in a (7:1):2 ratio.
-We treat the class with [100:][100:][ 100 : ] positive cases as head classes, $[30,100)$ cases as medium classes, and [:30)[:30)[ : 30 ) cases as tail classes.
-Consequently, from a total of 38,858 cases, the training set comprises 27,201 cases, and the test set includes 7,772 cases.
-Following on the dataset construction procedure, we perform classification task on two class sets at different granularities: (i) 5568 disorders + normal; (ii) 930 ICD-10-CM codes + normal.
-On different sets, we will have different head/medium/tail class set following our definition.
-Note that, the normal class is always treated as one of the head classes.
-Next, we will only talk about the abnormal classes.
-Disorders.
-At the disorder level, there are 85 head classes, 470 medium classes, and 5014 tail classes.
-Notably, as the case number of some tailed classes is extremely small, some classes may only appear in training or test sets, but not both.
-This issue only happens to the tail classes.
-It is important to note that our knowledge encoder enables to get embedding for unseen diseases, ensuring that evaluation is not affected by such challenge.
-As a result, there are 5,305/3,399 classes in our training and testing split, respectively.
-ICD-10-CM.
-At ICD-10-CM level, disorders with similar meanings have been mapped to same codes, for example, ‘tuberculosis’, ‘primary pulmonary tuberculosis’ all correspond to ‘A15 respiratory tuberculosis’.
-This merge operation results in more cases in one class, consequently more head classes.
-Specifically, we have 165 head classes, 230 medium classes, and 536 tail classes.
-Similarly, some tail classes may only be in training or testing split, resulting in 902/759 classes for our final training and testing splits, respectively.
-It is important to point that while we ensure the inclusion of head and medium classes in both the training and test sets, some tail classes are inevitably exclusive from certain divisions, which is treated as our future work to collect more samples of such disorders.
-We report the results separately for head/medium/tail classes, while focusing primarily on the head classes.
-5.1.2 Evaluation Metrics
-In this section, we describe the evaluation metrics in detail.
-Note that, the following metrics can all be calculated per class.
-For multi-class cases, we all use macro-average on classes to report the scores by default.
-For example, “AUC” for multi-class classification denotes the “Macro-averaged AUC on classes”.
-AUC.
-Area Under Curve [11] denotes the area under ROC (receiver operating characteristic) curve.
-This has been widely used in medical diagnosis, due to its clinical meanings and robustness in unbalanced distribution.
-AP.
-Average Precision (AP) is calculated as the weighted mean of precisions at each threshold.
-Specifically, for each class, we rank all samples according to the prediction score, then shift the threshold to a series of precision-recall points and draw the precision-recall (PR) curve, AP equals the area under the curve.
-This score measures whether the unhealthy samples are ranked higher than healthy ones.
-We report Mean Average Precision (mAP), which is the average of AP of each class.
-F1 and MCC.
-F1 score is the harmonic mean of the precision and recall.
-It is widely used in diagnosis tasks.
-MCC [18] denotes the Matthews Correlation Coefficient metric.
-It ranges from $-1$ to $1$ and can be calculated as follow:
-$\text{MCC}=\frac{\text{TN}\times\text{TP}-\text{FN}\times\text{FP}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}$
-Both metrics need a specific threshold to compute, and we choose one by maximizing F1 on the validation set following former papers [59, 79, 80, 84].
-Recall@FPR.
-We also report the recall scores at different false positive rate (FPR), i.e., sample points from class-wise ROC curves.
-Specifically, We report the Recall@0.01, Recall@0.05, Recall@0.1 score, denoting the recall scores at 0.01, 0.05, 0.1 FPR respectively.
-5.2 Transfer Learning to External Datasets
-In addition to evaluating on our own benchmark, we also consider transferring our final model to other external datasets, demonstrating its transferring abilities on image distribution shift and label space change.
-In the following sections, we start by introducing the used external datasets, that cover various medical imaging modalities and anatomies.
-Then we detail the fine-tuning settings.
-5.2.1 External Datasets
-We choose external evaluation dataset with the following principles:
-  • Imaging Modalities. We hope to cover most radiologic modalities in our external evaluation, e.g., 2D X-ray, 2D CT/MRI slices and 3D CT/MRI scans, demonstrating our model can benefit all of them;
-  • Human Anatomies. We hope to cover many human anatomies, e.g., brain, head and neck, chest, spine abdomen and limb, demonstrating our model can benefit all of them;
-  • Label Space.
-We hope to cover two cases, i.e., seen and unseen classes.
-Seen classes refer to those labels that have appeared in RP3D-DiagDS.
-Conversely, unseen extra classes denote those not included.
-As a result, we pick the following datasets, and report AUC scores to compare with others on these external evaluation.
-  • CXR14 [76] is a widely-used chest X-ray (2D) diagnosis dataset containing 112,120 frontal-view X-ray images of 30,805 (collected from the year 1992 to 2015) unique patients with 14 finding labels.
-We follow its official split and evaluate the SOTA [80] on the split.
-  • VinDr-Spine [60] is a spine X-ray (2D) diagnosis dataset comprising 10,469 images from 5,000 studies. We follow K-Diag [79] to use the 8 unique finding labels and the official split.
-  • VinDr-Mammo [60] is a mammography (2D) diagnosis dataset comprising 20,000 images (5,000 four-view scans).
-Each scan was manually annotated with a 5-level BI-RADS score.
-We view this as a multi-class classification task with the official split.
-  • ADNI [35] is a 3D brain MRI dataset focused on Alzheimer’s disease comprising 112141 images, including AD, MCI, CN and other classifications.
-We random split it follow 63,846/15,962 for training and testing and reproduce the state-of-the-art method on it.
-  • MosMedData [57] is a 3D chest CT dataset on 5-level COVID-19 grading comprising 1,110 images.
-We follow the official split.
-We random split it follow 888/222 for training and testing and reproduce the state-of-the-art method on it.
-5.2.2 Fine-tuning Diagnosis
-Our final model can serve as a pre-trained model and be fine-tuned on each downstream task to improve the final performance, demonstrating the merit of our dataset.
-Specifically, for the dataset with single image input, we simply adopt the pre-trained visual encoder module, i.e., discarding the fusion module.
-While for the datset with multi image input, we will use them all.
-In both cases, the final classification layer will be trained from scratch.
-In addition to using all available external training data, we also consider to use 1%, 10%, 30% portion data for few-shot learning.
-5.3 Implementation Details
-At training time, we consider two diagnose tasks in different granularities: Disorder-level classification (5569 classes) and ICD10-level classification (931 Classes).
-The image input will all be resized to $512\times 512\times D$ in height, width and depth respectively.
-For 3D scans, the depth is treated as a factor for ablation study, and will be discussed in experiment section.
-In vision encoder, we adopt two separated modules for normalization and a shared module to compute the final embedding.
-The detailed architecture design will also be further discussed in our ablation study.
-In fusion module, we use a 6-layer transformer encoder with learnable ‘[cls]’ token for final prediction.
-For augmentation, we adopt Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation, implemented from the MONAI [16] package.
-For optimization, we utilize the AdamW optimizer with a cosine learning rate curve, setting the maximum learning rate to $lr=1\times e^{-5}$, with an adjustable batch size ranging from 4 to 32 depending on the input image depth and model scale to avoid out-of-memory error.
-The total training duration spans 100 epochs, with the initial 5 epochs for warm-up.
-At fine-tuning stage, we adopt the similar model architecture and optimization setting.
-6 Results
-We conduct an ablation study to identify the optimal architecture and hyper-parameters for our model.
-Then, we present the evaluation results, focusing on the ICD-10-CM classification and disorders classification.
-Lastly, we use our dataset for large-scale pre-training, and finetune it on various external datasets, to demonstrate the model’s transferring ability.
-6.1 Ablation Study
-To explore the optimal model architecture and parameter configurations, including the fusion strategy, visual encoder architecture, the depth of 3D scans, and augmentation, we conduct a series of ablation studies on a subset of original dataset, comprising 200 disorder categories with most cases, termed as SubSet@200.
-In the default experiment setting, we use 16 as the 3D scan depth, and ResNet as visual backbone, without knowledge enhancement and augmentation strategy.
-While conducting ablation study on certain component, we keep other setting unchanged.
-6.1.1 Visual Encoder Architecture
-In this section, we investigate different backbone architectures for diagnosis, specifically, we compare the ResNet-based and the ViT-based models.
-As shown in Table 2, “Seperated” denotes the separated visual encoder to perform 2D and 3D normalization and “Shared” denotes the shared encoder part for both 2D and 3D scans.
-We make two observations, (i) ResNet-based architecture performs better than ViT-based ones, (ii) increasing the capacity of shared encoder is more beneficial, e.g., ResNet-34+ResNet-18 vs.
-ResNet-18+ResNet-34.
-(iii) Deeper ResNet structure has very limited improvement, e.g., ResNet-34+ResNet-34 vs.
-ResNet-18+ResNet-34.
+where the model is composed of a visual encoder $\Phi_{\text{visual}}$, a fusion module $\Phi_{\text{fuse}}$ and a classifier $\Phi_{\text{cls}}$.
+RP3D-DiagDS Statistics
+Here, we provide a detailed statistical analysis of our collected dataset, namely RP3D-DiagDS, from three aspects: modality coverage, anatomy coverage, and disease coverage.
+As shown in Supplementary Section E, compared with existing datasets, our dataset has great advantages in terms of anatomy richness, modal richness, and the number of images and categories.
+Furthermore, we split the dataset based on the number of cases on different levels.
+Analysis on Modality Coverage.
+RP3D-DiagDS comprises images from 9 modalities, namely, computed tomography (CT), magnetic resonance imaging (MRI), X-ray, Ultrasound, Fluoroscopy, Nuclear medicine, Mammography, DSA (angiography), and Barium Enema.
+Each case may include images from multiple modalities, to ensure the precise and comprehensive diagnosis of disorders.
+Overall, approximately 19.4% of the cases comprise images from two modalities, while around 2.9% involve images from three to five modalities.
+The remaining cases are associated with image scans from a single modality.
+The distribution of modalities among all abnormal and normal samples are illustrated in Figure 2a.
+Analysis on Anatomy Coverage.
+RP3D-DiagDS comprises images from various anatomical regions, including the head and neck, spine, chest, breast, abdomen and pelvis, upper limb, and lower limb, providing comprehensive coverage of the entire human body.
+The statistics are shown in Figure 2b.
+Figure 2: The data distribution analysis on RP3D-DiagDS.
+a The distribution of imaging modalities of abnormal (left) and normal (right) cases in RP3D-DiagDS.
+Each label is annotated with the class name, number of cases, and the corresponding proportion.
+b The distribution of imaging anatomies of abnormal (left) and normal (right) cases in RP3D-DiagDS.
+c Case distribution on image numbers.
+In the bar plot, We show the distribution for the number of images in one case.
+In RP3D-DiagDS, each case may include multiple images from patient history scans, different modalities, and different angles or conditions.
+d Case distribution on classes.
+We demonstrate the long-tailed distributions for disorder and ICD-10-CM classes.
+We also categorize these classes into three categories: “head class”, “body class” and “tail class” based on the number of cases.
+Notably, to better show the main part of the case distributions, we clip the axes, indicated by the dotted axes lines.
+Analysis on Disease Coverage.
+For both disorder and disease classification, each case can correspond to multiple disorders, resulting in RP3D-DiagDS a multi-label classification dataset.
+As shown in Figure 2d, the distributions exhibit extremely long-tailed patterns, rendering such 2D & 3D image classification problems as a long-tailed multi-label classification task.
+Totally there are 5568 disorders mapped into 930 ICD-10-CM classes.
+We define the “head class” category with case counts greater than 100, the “body class” category with case counts between 30 and 100, and the “tail class” category with case counts less than 30.
+Ablation Studies
+We first carry out a series of ablation studies to explore the optimal model architecture and parameter configurations, including the fusion strategy, visual encoder architecture, the depth of 3D scans, and augmentation.
+Note that, considering the original RP3D-DiagDS is too large to carry out extensive empirical experiments, we construct a subset by only keeping the top 200 classes with the most cases, accordingly, tracking the performance on these 200 classes for ablation studies.
+As simple experiment baselines, we use images with spatial resolution of 512x512, 16 as the 3D scan depth, and ResNet-18+ResNet-34 as the visual backbone, without knowledge enhancement or augmentation strategy, adopting the basic max pooling fusion strategy introduced in Supplementary Section B.
+While conducting an ablation study on certain component, we keep other settings unchanged.
+More results can be found in Supplementary Section C.
+Herein, we mainly describe the critical results on comparing image encoder architectures.
 Table 1: Ablation study on various settings of visual encoder architectures.
@@
 ResNet-34 ResNet-34 87.10 11.31 17.66 19.41 21.33 44.19 62.25
-6.1.2 Image Dimension
-Here, we aim to investigate the effect of input resolution, i.e., increasing the depth of input volumes.
-Due to the constraints of GPU physical memory, we experiment with 16, 24, and 32 as the depths for 3D input volume.
-We employ trilinear interpolation to resample 3D scans to the same size.
-Table 3 illustrates the performance by varying the depths of 3D images.
-It can be observed that an increase in depth can bring clear performance gain, suggesting that the detailed depth information is critical to perform diagnosis.
-Consequently, with more slices, the model tend to yield more favorable results.
-Table 3: Ablation study on different 3D scan depth size settings. Considering the computational cost, we evaluated the commonly seen 16, 24, 32 depth sizes on our datsets.
-3D Image Depth AUC AP F1 MCC R@0.01 R@0.05 R@0.1
-16 87.06 11.27 17.36 19.23 21.48 44.38 61.54
-24 87.49 11.41 17.84 19.39 21.76 45.72 62.53
-32 88.13 11.98 18.02 19.67 22.84 47.46 63.91
-6.1.3 Augmentation
-Here, we evaluate the effectiveness of augmentation on our dataset.
-Specifically, we adopt four augmentation strategies with $15\%$ probability each, namely, Gaussian Noise, Contrast Adjustment, Affine Variation, and Elastic Deformation.
-As a result, about half of the training data will be applied at least one augmentation in each training batch.
-As shown in Table 4, adopting data augmentation shows notable performance improvement.
-This improvement is particularly evident on AUC and AP when employing the ViT as the visual backbone while, still, ResNet-based model performs better.
-Table 4: Ablation on adopting augmentation strategy.
-In the table, we show the effectiveness of adopting reasonable data augmentation for our task.
-“ResNet” denotes the ResNet-based visual encoder architecture and “ViT” denotes the ViT-basde one.
-Augmentation Visual Encoder AUC AP F1 MCC R@F0.01 R@F0.05 R@F0.1
-YES ResNet 88.96 13.92 22.04 23.08 25.75 51.61 66.47
-ViT 84.22 8.69 16.36 17.18 19.83 38.70 54.04
-NO ResNet 87.06 11.27 17.36 19.23 21.48 44.38 61.54
-ViT 81.69 6.40 14.71 15.30 18.11 34.73 48.84
-6.1.4 Summary
-In conclusion, we find the ResNet-based model with augmentation strategy and unifying the 3D scan depth to 32 are the most suitable setting for our long-tailed case-level multi-modal diagnosis task, which will be used for training on the entire dataset in the following sections.
-6.2 Evaluation on Rad3D-DiagDS
-In this section, we train our model on the entire Rad3D-DiagDS training set, and evaluate on the test split respectively, as shown in Table 5, and the AUC curve comparison is shown in Figure 7.
-Specifically, we carry out the experiments at two levels, i.e., disorder and ICD-10-CM classes.
-We denote fusion module as FM for short and knowledge enhancement as KE.
-In cases where without FM or KE, we adopt max pooling on the predictions of different images from the same case (Check supplementary for more details), serving as a baseline.
-Then, we add the fusion module and knowledge enhancement step-by-step, to improve the model’s performance.
+ResNet-ViT ResNet-34 6-layer ViT 88.74 11.52 17.86 20.05 21.92 44.63 63.09
+ResNet-50 6-layer ViT 89.53 11.75 19.59 20.61 23.18 51.34 67.39
+ResNet-34 12-layer ViT 88.93 11.49 18.07 20.09 22.38 45.23 65.04
+ResNet-50 12-layer ViT 89.56 11.73 19.73 20.11 22.58 51.64 67.92
+Specifically, we investigate three backbone architectures variants for realizing unified diagnosis with arbitrary 2D and 3D scans, i.e., ResNet-based, ViT-based and ResNet-ViT-mixing.
+More details on architecture design can be found in Section 4.2.1.
+As shown in Table 1, we make three observations, (i) The performance of ResNet-ViT-mixing based model significantly surpasses that of ViT-based and ResNet-based ones.
+(ii) Enhancing the feature dimension of the vision encoder results in notable improvements, e.g.,ResNet-34 + 6-layerViT vs.
+ResNet-50 + 6-layerViT.
+(iii) Increasing the number of shared encoder layers in the ViT-based model does not yield a significant enhancement in performance but incurs substantial computational overhead, e.g., ResNet-50 + 6-layer ViT vs.
+ResNet-50 + 12-layer ViT.
+Based on the comprehensive ablation results shown in Supplementary Section C, we choose the ResNet-ViT-mixing model with augmentation strategy and unifying the 3D scan depth to 32 as the most suitable settings for our following long-tailed case-level multi-modal diagnosis ability evaluation.
+After deciding the basic configurations, we evaluate the diagnosis ability on both our proposed internal RP3D-DiagDS dataset and various external benchmarks.
+In internal evaluation, we mainly show the effectiveness of our newly proposed modules, considering the problem as a multi-label classification task under a long-tailed distribution.
+In external evaluation, we systematically investigate the boundary of our final model, serving as a diagnosis foundation model under both fine-tuning and zero-shot settings.
+Evaluation on Internal Benchmark
+In this part, our model is trained on the entire Rad3D-DiagDS training set (29536 cases), evaluated on its test split (7772 cases), as detailed in Table 2.
+The comparison of AUC curves is depicted in Figure 3.
+We conduct experiments at two levels—disorder and ICD-10-CM classes—to evaluate the model’s performance, both with and without text knowledge enhancement.
+These evaluations consider two integration strategies: early fusion and late fusion.
+For comprehensive definitions and applications of these fusion strategies, please see Supplementary Section B for more details.
+The model’s effectiveness is assessed using various metrics on the test set, including the AUC curve’s confidence interval (CI), to gauge its stability and consistency.
 Table 2: Classification results on Disorders and ICD-10-CM levels.
 In the table “FM” represents the fusion module and “KE” represents the knowledge enhancement strategy.
-We report the results on Head/Medium/Tail class sets separately.
-Granularity Class Methods Metrics
-FM
-KE
-AUC AP F1 MCC R@0.01 R@0.05 R@0.1
-Disorders Head
-91.91 13.56 22.08 23.66 30.88 60.50 74.70
-93.56 14.2 24.77 25.84 36.19 65.21 79.83
-94.24 15.33 25.68 26.71 37.06 66.55 81.37
-Medium
-93.62 9.19 17.16 20.42 28.24 59.32 73.80
-94.03 11.47 19.18 23.14 30.04 63.89 78.69
-94.69 12.38 20.64 24.07 31.52 65.34 78.73
-Tail
-88.55 4.34 8.03 12.96 9.13 27.46 41.43
-90.01 8.97 9.35 13.30 10.04 27.02 42.32
-90.64 9.25 9.89 14.38 10.98 27.98 43.53
-ICD-10-CM Head
-88.29 12.01 19.80 20.85 22.95 50.12 62.25
-89.55 13.89 21.92 23.41 25.04 52.19 67.26
-90.89 14.37 22.67 24.29 26.11 53.82 69.16
-Medium
-88.37 7.85 15.58 18.02 22.88 49.62 65.25
-89.90 8.98 17.33 19.89 24.37 51.33 64.93
-91.67 9.56 18.32 20.77 25.68 52.85 66.63
-Tail
-85.28 4.02 7.53 11.71 7.94 21.28 34.41
-86.34 4.77 8.81 12.85 8.83 22.36 37.11
-86.55 4.81 8.69 12.74 8.86 22.75 37.97
-(a) ROCs on Disorders (Head).
-(b) ROCs on Disorders (Medium).
-(c) ROCs on Disorders (Tail).
-(d) ROCs on ICD-10-CM (Head).
-(e) ROCs on ICD-10-CM (Medium).
-(f) ROCs on ICD-10-CM (Tail).
-Figure 7: ROC curves on Disorders and ICD-10-CM, including head/medium/tail parts respectively.
-The shadow in the figure shown the 95% CI (Confidence interval) and FM, KE are short for Fusion Module and Knowledge Enhancement.
-Discussion on Fusion Module.
-To demonstrate the effectiveness of fusion module, we start with a baseline without FM or KE.
-As shown in Table 5, adding the fusion module can greatly improve the results on Head, Medium and Tail classes at both disorder and ICD-10-CM levels, showing the critical role for case-level information fusion in diagnosis task.
-These results align well with our expectations.
-In clinical practice, the examination of one modality for a diagnosis is often insufficient.
+We report the results on Head/Medium/Tail category sets separately.
+Granularity Category Methods Metrics
+FM KE AUC AP F1 MCC R@0.01 F@0.05 R@0.1
+Disorders Head 92.44 15.34 24.95 26.14 34.11 63.70 76.70
+93.61 18.16 28.15 29.30 39.89 67.74 80.10
+94.41 20.27 30.21 32.27 41.93 71.09 81.15
+Medium 93.74 11.12 19.88 22.67 32.62 62.29 73.63
+94.79 13.51 22.96 25.80 37.63 66.25 78.06
+95.14 15.95 25.82 28.84 42.49 68.73 79.39
+Tail 88.10 6.04 10.35 15.06 9.52 22.48 32.94
+90.13 6.65 10.98 15.84 11.43 27.37 43.01
+90.96 7.75 12.68 17.49 13.13 28.88 44.08
+ICD-10-CM Head 89.83 12.31 20.27 21.58 23.81 52.05 68.45
+90.25 13.94 21.78 22.98 25.94 55.22 69.15
+91.27 14.59 22.81 25.12 27.87 57.75 72.32
+Medium 90.58 8.42 16.42 19.41 23.80 50.80 66.42
+91.12 9.18 18.02 20.05 27.10 53.77 67.45
+92.01 10.34 19.08 22.16 28.81 55.55 69.86
+Tail 86.91 4.82 8.64 12.70 10.13 29.82 51.04
+87.32 5.09 9.29 13.23 11.21 30.00 49.39
+88.11 5.57 10.48 14.68 12.75 30.11 51.77
+    *All indicators are presented as percentages, with the % omitted in the table.
+Analysis on Fusion Module (FM).
+To demonstrate the effectiveness of the fusion module, we start with a baseline that adopts max pooling on the predictions of different images from the same case (check Supplementary Section B for more details).
+Then, we add the fusion module and knowledge enhancement step-by-step, to improve the model’s performance.
+As shown in Table 2, adding the fusion module can greatly improve the results on Head, Medium, and Tail classes at both disorder and ICD-10-CM levels, showing the critical role of case-level information fusion in diagnosis tasks.
+These results align well with our expectations, as in clinical practice, the examination of one modality for a diagnosis is often insufficient.
 A thorough and meticulous diagnostic process typically involves an integrated review of all test results.
@@
 Our fusion model adeptly mirrors this comprehensive approach, demonstrating its effectiveness in simulating the nuanced process of clinical diagnosis.
-Discussion on Knowledge Enhancement.
-As shown in Table 5, incorporating knowledge enhancement further promotes the final diagnosis performance on both disorder and ICD-10-CM cases, validating the assumption that a better text embedding space trained by rich domain knowledge can greatly help the visual representation learning.
-Discussion on ROC Curves.
+Analysis on Knowledge Enhancement (KE).
+We experiment with the knowledge-enhanced training strategy, where we first leverage the rich domain knowledge to pre-train a knowledge encoder with natural language and use it to guide the visual representation learning for disease diagnosis.
+As shown in Table 2, knowledge enhancement further promotes the final diagnosis performance on both disorder and ICD-10-CM classifications, showing that a better text embedding space trained by rich domain knowledge can greatly help visual representation learning.
+Figure 3: ROC curves on Disorders and ICD-10-CM, including head/medium/tail parts respectively.
+The shadow in the figure shows the 95% CI (Confidence interval) and FM, KE is short for Fusion Module and Knowledge Enhancement.
+More ROC curves for each class are shown in Supplementary Section F
+Analysis on ROC Curves.
 The solid lines in the Figure 3 represents the median AUC value.
@@
 This procedure is repeated multiple times (1000), with the final curve representing the median of these values.
-The ROC curve is illustrated by thess solid lines.
-Accompanying each solid lines is shaded area, which denotes the 95% confidence interval (CI).
+The ROC curve is illustrated by the solid lines.
+Accompanying each solid line is a shaded area, which denotes the 95% confidence interval (CI).
 For each type of implementation, this shaded area is significantly higher than the portion above the solid line.
 This implies that the AUC value represented by the solid line is generally higher than the class-wise average across different data splits.
-As shown in the figure, the highest AUC value has reached 0.96, and the lowest value is 0.86, this observed pattern suggests, there exists a few categories that are more challenging to learn than others.
+As shown in the figure, the highest AUC value has reached 0.97, and the lowest value is 0.91, this observed pattern suggests, that there exists a few categories that are more challenging to learn than others.
+Analysis on Normal/Abnormal Diagnosis.
+In clinical scenarios, the proportion of normal cases is significantly higher compared to abnormal cases.
+Therefore, it is crucial to highlight the evaluation of the model’s ability to distinguish between normal and abnormal cases.
+As shown in Figure 5a, our model achieves an AUC of 82.61% for such diagnosis, given the significant disease variations in abnormal cases.
+Additionally, we also split out six test sub-sets based on the anatomy of cases, including Head and Neck, Chest, Abdomen and Pelvis, Upper Limb, Lower Limb, Spine, and Breast.
+In each sub-set, all cases are from the same anatomy, thus we can track the performance among a certain anatomy region.
+As shown in Figure 5a, the results on each anatomy are generally consistent, indicating that our model robustly demonstrates superior performance across different anatomical regions, without showing notable biases on any specific area.
+Analysis on Prediction Probability Distribution.
+To get a deeper and more intuitive assessment of the model’s capabilities, we analyzed its performance across different disease classes.
+Given the large number of diseases, we pick a total of six diseases, as shown in Figure 4, where we demonstrate the prediction probability distribution on cases in the test set for a certain class.
+All cases are categorized by their related anatomies, with positive cases corresponding to the considered disease class and negative cases divided into two types: intra-negative cases from the same anatomy, and inter-negative cases from other anatomies.
+Horizontal dashed lines in the bar plot denotes the threshold, with positive and negative cases from the same anatomy colored in red and blue, and negative cases from other anatomies in different shades of gray.
+Ideally, positive cases should have high scores, while both types of negative cases should have low scores.
+Therefore, the difference between the actual distribution and the ideal situation can effectively illustrate the model’s performance on the disease.
+As shown by the results, for the success classes, the model can distinguish both the intra-negative and inter-negative cases against the positive, while for the ordinary classes, the model may be confused with some intra-negative and positive cases.
+In the failure class, the model cannot pick out the positive cases.
+These results demonstrate that our internal evaluation requires the model to perform a two-stage differential diagnosis.
+The first step is to locate the possible diseases through the shooting anatomy, etc., and then capture the detailed lesion patterns features to make further diagnosis.
+Figure 4: The prediction probability distribution for different classes and saliency map visualization.
+a Based on the anatomies of each case, we split them into positive cases, intra-negative cases which are located in the same anatomy as the positive ones, and inter-negative cases which are located in other anatomies.
+The classification threshold score in the figure denotes the final comparison bar to transform the soft probabilities into binary true/false diagnosis results.
+The first three probability distribution figures depict the distributions of three relatively successful classes, where the model can clearly distinguish the inter-negative cases and the intra-negative cases are more confusing.
+We then show two ordinary classes.
+As shown by the distributions, most errors are caused by the intra-negative cases, and similarly, the inter-negative cases are easily dismissed as well.
+At last, we show a failure case where the model can hardly distinguish the positive and negative cases regardless of whether they are intra-negative or inter-negative.
+b Saliency map of the key frames.
+Red indicates the areas that the model focuses on when inferring the corresponding disease category.
+This indicates that RadDiag is capable of accurately identifying the locations of lesions or abnormal regions.
+Analysis on Saliency Map Visualization.
+Additionally, inspired by the proposal of explainable AI (XAI) [40], we utilize Score-CAM [74], a visual explanation method based on class activation maps, to visualize the areas that contribute more to make a decision on the target disease class.
+As shown in Figure 4b, we sample several cases from different modalities and anatomies to compute the saliency map.
+For instance, as depicted in Figure 4b(1) for “Ulnar styloid impaction syndrome”, our model can pinpoint the location of the ulna during inference for this category, without being distracted by other anomalies, such as rings worn on the four finger.
+In another case in Figure 4b(2), for the category of “Spondylodiscitis”, our model shows its ability to focus on the intervertebral disk, with heightened attention particularly on the region of abnormality.
+It proves our model’s precision in identifying the relevant anatomical structures for the disease categories, even amidst other potential distractions or anomalies within the same image, thus illustrating its robustness and specificity in medical disease diagnosis.
+Evaluation on External Benchmarks
+In this part, we further assess the transferring ability of our diagnosis model trained on RP3D-DiagDS.
+We adopt two settings to demonstrate the transferring ability to External Benchmarks of our model, i.e., zero-shot, and fine-tuning.
+First, in the zero-shot setting, we evaluate the effectiveness of our model for directly deployed in real clinical scenarios.
+In detail, as our dataset encompasses a wide range of disease categories affecting the entire body, it can be tested on most external datasets.
+More details on how we perform zero-shot transferring for our method and other baselines can be found in Section 4.3.3.
+We have adopted F1, MCC, and ACC for comparison.
+Second, our model can also serve as a pre-trained model and be fine-tuned on each downstream task to improve the final performance.
+Specifically, for the dataset with single image input, we simply adopt the pre-trained visual encoder module, i.e., discarding the fusion module.
+For the dataset with multi-image input, we inherit the entire model.
+In both cases, the final classification layer will be trained from scratch.
+In addition to using all available external training data, we also consider using 1%, 10%, and 30% portion data for few-shot learning.
+Figure 5: Analysis on normal/abnormal diagnosis.
+a The bar plot shows the AUC, AP, and Recall@0.05 of the Normal class on different body regions in the Internal test set of RP3D-DiagDS.
+b The radar figure shows the AUC and AP performance comparison between RadIN, BiomedCLIP, and Ours on external clinical datasets at five different anatomies.
+Zero-shot Results on Normal/Abnormal Diagnosis.
+In zero-shot assessment, we evaluate the transferring ability of our final model on normal/abnormal diagnosis for external dataset, e.g., Brain-Tumor (405 normal, 906 abnormal), POCUS (256 normal, 302 abnormal), CT-Kidney (1037 normal, 1453 abnormal), MURA (1667 normal, 1530 abnormal), and VinDr-Spine (1070 normal, 1007 abnormal), totaling 9634 cases.
+These datasets are derived from medical centers in different regions, thus is a good assessment of the model’s performance in distinguishing normal from abnormal cases under zero-shot conditions.
+We directly adopt the “normal” classifier head trained on our RP3D-DiagDS to obtain the prediction probability score and compare it with other foundation models, RadIN [53] and BiomedCLIP [82], in terms of AUC and AP on the five zero-shot datasets.
+As shown in Figure 5b, our model demonstrates significantly better results on all the five considered datasets, i.e., Brain-Tumor, POCUS, CT-Kidney, and MURA, compared to RadIN and BiomedCLIP.
+These findings show that our model consistently demonstrates superior diagnostic performance on zero-shot normal/abnormal diagnosis, across different anatomies under cross-center real clinical practices, significantly outperforming existing SOTA models.
+Zero-shot Results on Fine-grained Diagnosis.
+In addition to the normal/abnormal diagnosis, as shown in Table 3, we also conduct zero-shot evaluation on fine-grained diagnosis tasks introduced by each representative external datasets, spanning six anatomies, including the head and neck, chest, breast, abdomen and pelvis, and spine.
+These datasets cover five modalities, i.e., CT, MRI, X-ray, Ultrasound, and Mammography, allowing for a comprehensive evaluation of our model’s capabilities.
+As baselines, we have included RadIN [53] and BiomedCLIP [82] for comparison in the zero-shot setting.
+The results demonstrate that our model outperforms the other two models in overall performance, indicating strong transfer learning capabilities.
+Interestingly, our model’s zero-shot performance can sometimes be even comparable to the results obtained by the model trained from scratch on a subset of the corresponding dataset, demonstrating the superiority of our RadDiag for diagnostic capabilities, which cannot be achieved by former foundation models.
+Table 3: Zero-shot results on 6 external datasets.
+For our model, we avoid exposing to these external data for zero-shot experiments and use the threshold corresponding to the highest F1 score on the evaluation set.
+For RadIN and BiomedCLIP, we use the threshold corresponding to the highest F1 score on each individual external dataset.
+We also report the supervised training results with partially down-stream task external data, denoting as “S.T.”, for demonstrating what diagnosis performance our model can achieve zero-shotly.
+We have evaluated F1, MCC and ACC, and all the results reported in the table are the averaged values across all classes.
+In the table, "US" represents ultrasound and "MG" represents mammography.
+Dataset Anatomy Modality Dim Method External Data F1 $\uparrow$ MCC $\uparrow$ ACC $\uparrow$
+Brain-Tumor Head&Neck MRI 2D RadIN Zero-shot 22.29 3.23 63.63
+BiomedCLIP 56.25 32.81 57.21
+Ours 61.09 47.37 73.38
+S.T. n=571(10%) 61.77 46.69 72.17
+POCUS Chest US Video RadIN Zero-shot 2.22 0.17 66.67
+BiomedCLIP 50.33 11.73 45.82
+Ours 65.02 47.97 79.84
+S.T. n=223(1%) 64.20 48.31 80.58
+CT-KIDNEY Abdom&Pelvis CT 2D RadIN Zero-shot 38.1 17.57 64.75
+BiomedCLIP 49.38 29.58 72.21
+Ours 56.50 43.18 69.96
+S.T. n=995(10%) 55.73 42.28 66.95
+MURA Limb X-ray 2D RadIN Zero-shot 31.26 4.37 52.94
+BiomedCLIP 68.36 14.96 54.70
+Ours 67.64 40.15 77.51
+S.T. n=3680(10%) 66.71 41.98 78.32
+VinDr-Spine Spine X-ray 2D RadIN Zero-shot 4.34 0.91 87.49
+BiomedCLIP 26.49 12.79 61.3
+Ours 25.93 19.40 83.07
+S.T. n=84(1%) 24.73 20.61 84.46
+TCGA Chest&Breast CT&MG 3D RadIN Zero-shot - - -
+BiomedCLIP - - -
+Ours 72.18 57.20 83.47
+S.T. n=173(30%) 70.33 52.26 77.41
+    All indicators are presented as percentages, with the % omitted in the table.
+Finetuning Results.
+As shown in Table 4, while finetuning our model on external datasets, we see significant performance improvement on all 22 external datasets with different data portions, compared to models trained from scratch.
+In most cases, our model can even surpass the specialist SOTAs which are designed carefully for the targeting task, demonstrating that publicly shared medical data on the Internet is a tremendous and valuable resource that can serve as a superior large-scale supervised training dataset for the medical domain.
+Table 4: The AUC score comparison on various external datasets.
+For each dataset, we experimented with different training data portions, denoted as 1% to 100% in the table.
+For example, 30% represents we use 30% of data in the downstream training set for finetuning our model or training from scratch.
+“SOTA” denotes the best performance of former works (pointed with corresponding reference) on the datasets.
+We mark the gap between ours and training from scratch on the subscript of $\uparrow$ in the table.
+Dataset 1% 10% 30% 100% SOTA
+Scratch Ours Scratch Ours Scratch Ours Scratch Ours
 […]
-    The numbers are borrowed from the referred papers. We are strictly align with them in train and test split.
-6.3 External Evaluation
-In this section, we explore the transferring ability of our final model, where we finetune the model on different downstream datasets across various imaging modalities, anatomies and target classes.
-As shown in Table 6.2, we can see significant performance improvement on various datasets with different data portions, compring to models trained from scratch.
-Additionally, in most cases, our model can also surpass former SOTAs significantly without any task-specific designs, e.g., architectures, loss functions, demonstrating that our RP3D-DiagDS can also serve as a superior large-scale supervised pre-training dataset for medical domain.
-7 Clinical Impact
-In this paper, we consider the problem of multi-modality, multi-anatomy, long-tailed disease diagnosis, which has great practical clinical meanings.
-First, in clinical practice, patients may get multiple radiologic examinations during their long-term treatment progress from various medical departments.
-However, existing work on disease classification accepting one image scan, can hardly handle such circumstance, hindering the development of accurate and comprehensive diagnostic models.
-Our work compensate this and better meet the practical scenarios.
-Second, comparing with existing diagnosis works that focus on common general disease classes, our efforts on long-tailed rare classes are more vital for clinical usage.
-For common diseases, AI diagnosis system, usually, can only help accelerate diagnosis procedure for clinicians, while the hint on rare classes is also critical.
-Third, our final model can support finetuning on external diagnosis task.
-In clinical application, local medical centers can leverage this superiority even with a few training samples.
-This is remarkable especially considering that, for rare diseases, only a very few cases could be accessible in reality.
-8 Limitation
-Despite the effectiveness of our proposed dataset, and architecture, there remains some improvements: First, on model design, in fusion step, we can use more image tokens to represent a scan rather than a pooled single vector.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+    The numbers are borrowed from the referred papers.
+  • The four SOTAs [54, 44, 81, 13] are relatively old. Thus, their network architecture is weaker than the classical 3D ResNet trained from scratch.
+    All indicators are presented as percentages, with the % omitted in the table.
+3 DISCUSSION
+There have been a number of open-source datasets for disease diagnosis.
+Notably, large-scale datasets such as NIH ChestX-ray [75], MIMIC-CXR [39], and CheXpert [34] stand out for their comprehensive collection of annotated X-ray images, facilitating extensive research and advancements in automated disease classification.
+However, it’s important to note three key limitations.
+First, these open-source diagnosis datasets mainly consist of chest X-rays [59, 15, 75, 34, 51, 30, 37, 69, 26, 55, 39, 21], thus are 2D images.
+There are only a few 3D datasets available, with a limited number of volumes [3, 2, 4, 42, 52, 32, 12, 8, 11].
+Second, a significant portion of these large-scale datasets focuses solely on binary classification of specific diseases [69, 26, 30, 37].
+Third, existing large-scale datasets contain a broad range of disease categories in different granularities [15, 53].
+For example, in PadChest [15], disorders exist at different levels, including infiltrates, interstitial patterns, and reticular interstitial patterns.
+Such variation in granularity poses a great challenge for representation learning.
+In summary, these datasets fall short of meeting real-world clinical needs, which often involve complex, multimodal, and multi-image data from a single patient.
+Therefore, constructing a dataset that mirrors the intricacies of actual clinical scenarios is necessary.
+In comparison, the large-scale radiology dataset we construct covers a large number of categories of diseases, including various common modalities.
+The prevailing paradigm in earlier diagnostic models is a specialized model trained on a limited range of disease categories.
+These models focus on specific imaging modalities and are targeted towards particular anatomical regions.
+Specifically, ConvNets are widely used in medical image classification due to their outstanding performance, for example, [23, 71, 21] have demonstrated excellent results on identifying a wide range of diseases.
+Recently, Vision Transformer (ViT) has garnered immense interest in the medical imaging community, numerous innovative approaches  [22, 78, 61, 83, 77, 49, 28] have emerged, leveraging ViTs as a foundation for further advancements in this field.
+The other stream of work is generalist medical foundation models  [73, 79, 84].
+These models represent a paradigm shift in medical AI, aiming to create versatile, comprehensive AI systems capable of handling a wide range of tasks across different medical modalities, by leveraging large-scale, diverse medical data.
+MedPaLM M [73] reaches performance competitive with or exceeding the state-of-the-art (SOTA) on various medical benchmarks, demonstrating the potential of the generalist foundation model in disease diagnosis and beyond.
+RadFM [79] is the first medical foundation model capable of processing 3D multi-image inputs, demonstrating the versatility and adaptability of generalist models in processing complex imaging data.
+However, the development of GMAI models requires substantial computational power, which is often impractical for exploring sophisticated algorithms in academic labs.
+In contrast, leveraging our proposed RP3D-DiagDS dataset, our final model RadDiag is capable of performing universal disease diagnosis and classification tasks within a constrained parameter scope, transcending the limitations of focusing on specific anatomies or disease categories.
+This capability offers a more versatile and efficient approach to healthcare management, resulting in significant clinical and research implications compared with precious AI models trained on former diagnostic datasets as discussed in Supplementary Section E.
+Our impacts and contributions can be summarized in the following three aspects:
+Model Development for Case-Level Diagnosis.
+The RP3D-DiagDS dataset facilitates the creation of AI models capable of case-level diagnosis utilizing multi-modal and multi-scan inputs.
+This is a critical advancement since patients often undergo multiple radiological examinations across different medical departments during their treatment.
+Traditional disease classification models, which typically process single-image scans, are inadequate for such complex scenarios.
+Our dataset, with its case-level labels, promotes the development of comprehensive diagnostic AI systems, thereby addressing a significant gap in clinical practice.
+Long-Tailed Rare Disease Diagnosis.
+Unlike existing datasets that primarily cover common diseases, the RP3D-DiagDS encompasses over 5,000 disorders, including a significant number of rare diseases.
+This focus on long-tailed rare disease diagnosis is particularly important for clinical applications, as AI systems can readily assist in diagnosing common diseases but often fall short in identifying rare conditions.
+Our dataset thus plays a crucial role in enhancing the diagnostic process for rare diseases, which are typically more challenging to diagnose due to their lower prevalence.
+Value of Online Radiology Resources.
+The RP3D-DiagDS dataset, collected from the Internet, serves as a valuable resource for training disease diagnosis model, especially for local medical centers dealing with rare diseases.
+Given the scarcity of cases for such diseases, our dataset’s ability to support few-shot learning scenarios is of paramount importance.
+It demonstrates the value that, such publicly shared medical data on the Internet represents a colossal and high-quality resource, holding the potential to pave the way for constructing generalist models aimed at advancing AI4healthcare.
+Despite the effectiveness of our proposed dataset, and architecture, there remains certain limitations: First, on model design, in the fusion step, we can use more image tokens to represent a scan rather than a pooled single vector.
 The latter may lead to excessive loss of image information during fusion.
-The model size can be further increased to investigate the effect of model capacities; Second, new loss functions should be explored to tackle such large-scale, long-tailed disorder/disease classification task.
-Third, on disorder to ICD-10-CM mapping progress, the annotators are labeled on class-name level, i.e., only disorder names are provided, causing some ambiguous classes unable to find strict corresponding ICD-10-CM codes.
+The model size can be further increased to investigate the effect of model capacities; Second, new loss functions should be explored to tackle such large-scale, long-tailed disorder/disease classification tasks.
+Third, on the disorder to ICD-10-CM mapping progress, the annotators are labeled on class-name level, i.e., only disorder names are provided, causing some ambiguous classes unable to find strict corresponding ICD-10-CM codes.
 Though we mark out this class in our shared data files, if providing more case-level information, the mapping could be more accurate.
-We treat these as future work.
+We treat this as future work.
+4 METHODS
+In this section, we will detail the following aspects separately: dataset construction, model design, and experiment implementation.
+4.1 Dataset Construction
+In this section, we present the details of our dataset, RP3D-DiagDS, that follows a similar procedure as RP3D [79].
+Specifically, cases in our dataset are sourced from the Radiopaedia website [5] – a growing peer-reviewed educational radiology resource website, that allows clinicians to upload 3D volumes to better reflect real clinical scenarios.
+Additionally, all privacy issues have already been resolved by the clinicians at uploading time.
+It is worth noting that, unlike RP3D [79] that contains paired free-form text description and radiology scans for visual-language representation learning, here, we focus on multi-modal, multi-anatomy, and multi-label disease diagnosis (classification) under extremely unbalanced distribution.
+Overall, the proposed dataset contains 40,936 cases, of 195,010 images from 9 diverse imaging modalities and 7 human anatomy regions, note that, each case may contain images of multiple scans.
+The data covers 5,568 different disorders, that have been manually mapped into 930 ICD-10-CM [1] codes.
+We describe the procedure of our data curation process, shown in Figure 6.
+Specifically, we collect three main components from each case on Radiopaedia webpage, namely, “Patient data”, “Radiology images”, and “Articles”  (example webpage shown in Figure 6).
+“Patient data” includes brief information about this patient, for example, age, gender, etc.
+“Radiology images” denote a series of radiology examination scans.
+Notably, when uploading the cases, the physicians have also picked a default key slice to support diagnosis for each scan in “Radiology images”, we also crawl this as a piece of extra information.
+Figure 2a provides a statistical analysis on the number of images within one single case.
+“Articles” contains links to related articles named with corresponding disorders, which are treated as diagnosis labels and have been meticulously peer-reviewed by experts in Radiopaedia Editorial Board [^1] https://radiopaedia.org/editors.
+Figure 6: The data collection pipeline.
+First, we crawl data from an open-sourced and expert-checked website, including case descriptions, 2D/3D radiology images, and related Radiopaedia articles.
+Then, we filter article names by large language model tools (GPT-4) and check manually if mismatch.
+Finally, we map the article names into ICD-10-CM codes, cross-checked by a 10-year clinician.
+To start with, we collect 50,970 cases linked with 10,670 articles from the Radiopaedia website.
+However, naively adopting the article titles as disorders may lead to ambiguities from two aspects: (i) not all articles are related to disorders, (ii) article titles can be written with different granularity, for example, “pneumonia” and “bacterial pneumonia” can be referred to as different disorders, though they should ideally be arranged in a hierarchical structure, (iii) normal cases from Radiopaedia are not balanced in modalities and anatomies.
+Next, we discuss the 3-stage procedure proposed to alleviate the above-mentioned challenges.
+Article Filtering.
+We leverage the GPT-4 [60] to automatically filter the article list, keeping those referring to disorders.
+Specifically, taking inspiration from the self-consistency prompts [76], we design two different query prompts with similar meanings, as shown in Figure 6.
+An article name is labeled as a disorder if GPT-4 consistently gives positive results from both prompts, while for those GPT-4 gives inconsistent results, we manually check them.
+To measure the quality of filtering, we randomly sample a portion of data for manual checking to control its quality.
+The confusion metrics are shown in Figure 6.
+The $100\%$ precision score indicates our auto-filtering strategy, can strictly ensure the left ones to be disordered.
+Eventually, 5342 articles can pass the first auto-criterion and 226 pass the second manually checking, resulting in 5,568 disorder classes, as shown in Figure 6b.
+Cases not linked to any articles are excluded, ultimately yielding 38,858 cases.
+Mapping Disorders to ICD-10-CM.
+Upon getting disorder classes, they may fall into varying hierarchical granularity levels, here, we hope to map them to internationally recognized standards.
+For this purpose, we utilize the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes, which is a customized version of the ICD-10 used for coding diagnoses in the U.S.
+healthcare system.
+The ICD was originally designed as a healthcare classification system, providing a system of diagnostic codes for classifying diseases, including nuanced classifications of a wide variety of signs, symptoms, abnormal findings, complaints, social circumstances, and external causes of injury or disease.
+Specifically, we hire ten medical Ph.D.
+students to manually map the article titles into ICD-10-CM codes.
+After cross-checking by a ten-year clinician, we have mapped the 5,568 disorders into the corresponding ICD-10-CM codes.
+We unify various disorders into the first hierarchical level of ICD-10-CM code tree, resulting in 930 ICD-10-CM classes.
+For example, both “J12.0 Adenoviral pneumonia” and “J12 Viral pneumonia” are mapped to the code “J12 Viral pneumonia, not elsewhere classified”, dismissing the ambiguities caused by diagnosis granularity.
+Note that, we will provide the ICD-10-CM codes along with the original 5,568 classes, each accompanied by its corresponding definition, thus, the dataset can be employed for training diagnosis or visual-language models.
+Adding Normal Cases.
+Despite there exist normal cases on Radiopaedia [^2] https://radiopaedia.org/articles/normal-imaging-examples, covering most anatomies and modalities, we additionally collect more normal cases from additional external datasets, including Brain-Tumor-17, CheXpert, Vertebrase-Xray, KneeMRI, and VinDr-Mammo, and MISTR [^3] https://mistr.usask.ca/odin/, with images available for research under a CC-BY-NC-SA 4.0 license, i.e., simply adding the normal case part introduced in them.
+The expanded normal cases include 2503 cases.
+Finally, we get 40,936 cases containing 195,010 images labeled by 5,568 disorder classes and 930 ICD-10-CM classes and will continually maintain the dataset, growing the case number.
+4.2 Model Design
+Here, we aim to initiate a preliminary investigation on computational architectures for large-scale, long-tailed disease diagnosis on radiology, specifically, we have defined the problem of case-level multi-label classification, and elaborate the architecture in Sec.
+4.2.1, and knowledge-enhanced training strategy in Sec.
+4.2.2.
+4.2.1 Architecture
+Figure 7: The overview of our method.
+Three parts demonstrate our proposed visual encoders and fusion module, together with the knowledge enhancement strategy respectively.
+a The three types of vision encoder, i.e., ResNet-based, ViT-based , and ResNet-ViT-mixing.
+b The architecture of the fusion module.
+The figure shows the transformer-based fusion module, enabling case-level information fusion.
+c The knowledge enhancement strategy.
+We first pre-train a text encoder with extra medical knowledge with contrastive learning, leveraging synonyms, descriptions, and hierarchy.
+Then we view the text embedding as a natural classifier to guide the diagnosis classification.
+Our proposed RadDiag architecture consists of two key components: (i) a visual encoder that processes 2D or 3D input scan; (ii) a transformer-based fusion module, merging all information to perform case-level diagnosis.
+Visual Encoder. We consider two popular variants of the visual encoder, namely, ResNet [29] and ViT [24], as shown in Figure 7a. The visual encoding progress can generally be formulated as:
+$v_{i}=\Phi_{\text{visual}}(x_{i})\in\mathcal{R}^{d}.$ (2)
+where $x_{i}\in\mathcal{R}^{C\times H\times W\times(D)}$ denotes the input scan, $C,H,W$ refer to the image channel, height, and width of the input scan respectively.
+$D$ is optional, and only available for 3D input scans.
+The main challenge for architecture design comes from the requirement to process scans in both 2D and 3D formats.
+Here, we train separate normalization modules to convert the 2D or 3D inputs into feature maps of the same resolution and further passed into a shared encoding module.
+We here adopt three visual encoding variants to handle the challenge:
+  • ResNet-based.
+For 3D scans, they are first fed into 3D ResNet, followed by average pooling on depth to aggregate the information of the extra dimension; while for 2D scans, they are fed to the 2D ResNet to perform the same down-sampling ratio on height and width as for 3D.
+After normalization, both 2D and 3D scans are transformed into feature maps with the same resolution, $f_{i}\in\mathcal{R}^{d_{\text{res}}\times h\times w}$, where $d_{\text{res}}$ is the intermediate feature dimension and $h,w$ denote the normalized size.
+Then, the feature maps are passed into a shared ResNet to get the final visual embedding.
+  • ViT-based.
+For 3D scans, we convert the input volume into a series of non-overlapped 3D cubes, and pass them into MLP projection layers, to get vector embeddings; while for 2D scans, the input scan is broken into 2D patches, and projected into vector embeddings with another set of MLP layers.
+As ViT enables handling sequences of variable tokens, we can now pass the resulting vectors into a shared ViT to get the final visual embedding by confusing the feature embedding from arbitrary patch tokens.
+For position embedding, we adopt two sets of learnable position embeddings, one for 2D input, and the other for 3D, further indicating the network about the input dimension.
+  • ResNet-ViT-mixing.
+Inspired by former works on pseudo-3D residual networks [65].
+We, herein, also design a new 2D + 1D network architecture mixing both ResNet and ViT structures to encode 3D or 2D scans consistently.
+Specifically, for 3D scans, we treat them as a sequence of 2D slices, and for 2D scans, we also treat them as a sequence of 2D slices of length 1.
+We first adopt 2D ResNet to get normalized feature embedding per slice as $f_{i}\in\mathcal{R}^{d_{\text{res}}\times h\times w}$.
+We then use transformer architecture to perform aggregation on the feature of slice tokens, and thus is similar to ViT-based method.
+Thanks to the flexibility of transformer architecture, the model can handle arbitrary number of tokens, obtaining the final visual embedding by integrating features from varying slice feature tokens.
+Fusion Module.
+For case-level diagnosis, we propose to aggregate information from multiple scans with a trainable module.
+As shown in Figure 7b, we adopt transformer encoders, specifically, we first initialize a set of learnable modality embeddings, denoted as $\{p_{1},\dots,p_{M}\}$, where $M$ denotes the total number of possible imaging modalities.
+Given certain visual embedding ($v_{i}$) from modality $j$, we first add the corresponding modality embedding ($p_{j}$) to it, indicating which radiologic modality it is from.
+Then, we feed all visual embeddings from one case into the fusion module, and output the fused visual embedding $v_{\text{fuse}}\in\mathcal{R}^{d}$, from the “[CLS]” token, similar to paper [24], denoted as:
+$v_{\text{fuse}}=\Phi_{\text{fuse}}(v_{1}+p_{m_{1}},v_{2}+p_{m_{2}},\cdots,v_{S}+p_{m_{S}})\in\mathcal{R}^{d},$ (3)
+where $v_{i}$ denotes the visual embedding for each image scan and $m_{i}$ denotes the its corresponding imaging modality, like MRI or CT.
+Till here, we have computed the case-level visual embedding, which can be passed into a classifier for disease diagnosis.
+We adopt a knowledge-enhanced training strategy [77, 83], that has shown to be superior in long-tailed recognition problems, as detailed in the following section.
+4.2.2 Knowledge-enhanced Training
+With knowledge enhancement, we hope to leverage the rich medical domain knowledge to enhance the long-tailed classification.
+Our key insight is that the long-tailed diseases may fundamentally have some shared symptoms or radiological pathologies with the head classes, which could be encoded in text format.
+In detail, we first pre-train a text encoder with medical knowledge, termed as knowledge encoder, where names of similar disorders or diseases are projected to similar embeddings, for example, “lung disease” is closer to “pneumonia” in the embedding space than “brain disease”.
+Then, we freeze the text embeddings, and use them to guide the training of the vision encoder, as shown in Figure 7c.
+Knowledge Encoder Pre-training.
+We leverage several knowledge bases to pre-train a knowledge encoder, including Radiopaedia, ICD10-CM, and UMLS.
+Specifically, for each disorder term, we collect its definitions, radiologic features from Radiopaedia articles, synonyms, clinical information, and hierarchy structure from ICD10-CM, as well as definitions from UMLS.
+We aim to train the text encoder with the three considerations, i.e., synonyms, hierarchy, and descriptions.
+First, if two terms are identified as similar synonyms, we expect them to be close in text embedding space, like “Salmonella enteritis” and “Salmonella gastroenteritis”.
+Second, in the context of hierarchical relationships, we expect that if a disease is a fine-grained classification of another disease, its embeddings should be closer than those of unrelated diseases.
+For example, “J93, Pneumothorax and air leak” should exhibit closer embeddings with “J93.0, Spontaneous tension pneumothorax” than with “J96.0, Acute respiratory failure”.
+Third, for terms associated with descriptions, radiological features, or clinical information, we expect their embeddings to be close to the text embedding space.
+For example, “Intestinal Neuroendocrine Tumor” and “A well-differentiated, low or intermediate grade tumor with neuroendocrine differentiation that arises from the small or large intestine”.
+We start from an off-the-shelf text encoder, namely, MedCPT-Query-Encoder [38], and adopt contrastive learning for further finetuning.
+Given a target medical terminology name encoded by the text encoder, denoted as $f_{\text{tar}}$, its corresponding medical texts, i.e., synonyms, containing terminologies or related descriptions are treated as positive cases.
+Similarly, we encode them with the text encoder, denoted as $f^{+}$, and other non-related text embedding is treated as negative cases $f^{-}$.
+Note that, we keep positive and negative cases consistent in format, e.g., when the positive case is related to a description sentence, the negative cases are always some non-related description sentences instead of the short name words.
+The final objective can be formulated as:
+$\mathcal{L}_{\text{knowledge}}=-\log\frac{e^{f_{\text{tar}}^{T}\cdot f^{+}/\tau}}{\sum_{n=1}^{N}e^{f_{\text{tar}}^{T}\cdot f^{-}_{n}/\tau}},$ (4)
+where $\tau$ refers to the temperature, and $N$ denotes total sampled negative cases.
+By optimizing the contrastive loss we can further finetune the text encoder, resulting in a knowledge encoder, termed as $\Phi_{k}$.
+Knowledge-guided Classification.
+After training the knowledge encoder, we use it to encode the disorder/disease names into text embeddings, for example, denoting the names as $\{T_{1},T_{2},\dots,T_{c}\}$, where $T_{j}$ is a disorder/disease name like “pneumonia” or “lung tumor”.
+We embed these free texts with the knowledge encoder as:
+$t_{j}=\Phi_{k}(T_{j})\in\mathcal{R}^{d},$ (5)
+where $t_{j}$ denotes the text embeddings that are used for computing dot product with visual embeddings for case-level diagnosis:
+$p=v_{\text{fuse}}\cdot t\in\mathcal{R}^{c},$ (6)
+where $p$ is the final result. We use classical binary cross entropy (BCE) loss as the final training objective, denoted as:
+$\mathcal{L}=-\sum_{i=1}^{c}\mathcal{Y}_{i}\cdot\log(p_{i})+(1-\mathcal{Y}_{i})\cdot\log(1-p_{i}).$ (7)
+4.3 Evaluation Details
+In this section, we will describe our evaluation details.
+First, we build up an internal test-set based on our proposed RP3D-DiagDS, serving for assessing the models’ ability on case-level, multi-modal, multi-scan, long-tailed disease diagnosis.
+Then we introduce more external datasets for assessing the transferring ability of our final model on other clinical benchmarks.
+Lastly, we introduce the used evaluaiton metrics, along with the compared baselines.
+4.3.1 The Internal Test-Set
+To start with, for better assessing the model’s ability on case-level diagnosis, we split a test set from our own proposed datasets, RP3D-DiagDS to perform internal evaluation.
+Specifically, we randomly split our dataset into training (train and validation) and testing sets in a (7:1):2 ratio .
+We treat the class with [100:][100:][ 100 : ] positive cases as head classes, $[30,100)$ cases as medium classes, and [:30)[:30)[ : 30 ) cases as tail classes.
+Consequently, from a total of 40,936 cases in our proposed RP3D-DiagDS, the training set comprises 29,536 cases, and the test set includes 7,772 cases.
+Following on the dataset construction procedure, we perform classification tasks on two class sets at different granularities: (i) 5568 disorders + normal; (ii) 930 ICD-10-CM codes + normal.
+On different sets, we will have different head/medium/tail class sets following our definition.
+Note that, the normal class is always treated as one of the head classes.
+Next, we will only discuss the abnormal classes.
+Disorders.
+At the disorder level, there are 85 head classes, 470 medium classes, and 5014 tail classes.
+Notably, as the case number of some tailed classes is extremely small, some classes may only appear in training or test sets, but not both.
+This issue only happens to the tail classes, while our knowledge encoder (as detailed in Sec.
+4.2.2) enables embedding for unseen diseases, ensuring that evaluation is not affected by such challenge.
+As a result, there are 5,260/3,391 classes in our training and testing split, respectively.
+ICD-10-CM.
+At the ICD-10-CM level, disorders with similar meanings have been mapped to the same codes, for example, “tuberculosis”, and “primary pulmonary tuberculosis” all correspond to “A15 respiratory tuberculosis”.
+This merge operation results in more cases in one class and, consequently more head classes.
+Specifically, we have 165 head classes, 230 medium classes, and 536 tail classes.
+Similarly, some tail classes may only be in training or testing split, resulting in 894/754 classes for our final training and testing splits, respectively.
+It is important to point out that while we ensure the inclusion of head and medium classes in both the training and test sets, some tail classes are inevitably exclusive from certain divisions, which is treated as our future work to collect more samples of such disorders.
+We report the results separately for head/medium/tail classes while focusing primarily on the head classes.
+4.3.2 The External Test-Sets
+In addition to evaluating on our own testset, we also consider zero-shot or finetuning evaluation on other external datasets, demonstrating its transferring abilities on image distribution shift and label space change.
+In the following sections, we start by introducing the used external datasets, that cover various medical imaging modalities and anatomies.
+Then we detail the fine-tuning settings.
+We choose an external evaluation dataset with the following principles:
+  • Imaging Modalities. We hope to cover most radiologic modalities in our external evaluation, e.g., 2D X-ray, 2D CT/MRI slices, and 3D CT/MRI scans demonstrating our model can benefit all of them;
+  • Human Anatomies. We hope to cover many human anatomies, e.g., brain, head and neck, chest, spine abdomen, and limb, demonstrating our model can benefit all of them;
+  • Label Space.
+We hope to cover two cases, i.e., seen and unseen classes.
+Seen classes refer to those labels that have appeared in RP3D-DiagDS.
+Conversely, unseen extra classes denote those not included.
+As a result, we pick the following datasets and report AUC scores to compare with others on the external evaluation.
+  • CXR14 [75] is a widely-used chest X-ray (2D) diagnosis dataset containing 112,120 frontal-view X-ray images of 30,805 (collected from the year 1992 to 2015) unique patients with 14 finding labels.
+We follow its official split and evaluate the SOTA [78] on the split.
+  • VinDr-Spine [56] is a spine X-ray (2D) diagnosis dataset comprising 10,469 images from 5,000 studies. We follow K-Diag [77] to use the 8 unique finding labels and the official split.
+  • VinDr-Mammo [56] is a mammography (2D) diagnosis dataset comprising 20,000 images (5,000 four-view scans).
+Each scan was manually annotated with a 5-level BI-RADS score.
+We view this as a multi-class classification task with the official split.
+  • ADNI [36] is a 3D brain MRI dataset focused on Alzheimer’s disease comprising 112141 images, including AD, MCI, CN and other classifications.
+We randomly split it followed 63,846/15,962 for training and testing and reproduced the state-of-the-art method on it.
+  • MosMedData [54] is a 3D chest CT dataset on 5-level COVID-19 grading comprising 1,110 images.
+We follow the official split.
+We randomly split it following 888/222 for training and testing and reproduced the state-of-the-art method on it.
+  • NSCLC (Radiogenomics) [9] is a 3D chest dataset containing CT and PET scans.
+There are a total of 211 subjects with Non-small Cell Lung Cancer.
+We split the dataset based on the overall cancer stage (4 classes).
+The training and testing set are randomly split as 169 and 42.
+  • TCGA (Merged) is a 3D chest dataset composed of three chest-related data sets: TCGA-LUSC [41], TCGA-LUAD [7] and TCGA-BRCA [48].
+It contains three different cancers and a total of 356 cases, which is randomly split into 285 for training and 71 for testing.
+  • ISPY1 [33] is a breast 3D MRI dataset from patients in the I-SPY 1/ACRIN 6657 trials.
+It contains a total of 222 subjects with Breast Cancer, and we selected 755 samples among them as the data set.
+We randomly split the 604 for training and 151 for training.
+  • ChestX-Det10 [50] is a 2D chest X-ray subset of NIH ChestX-14. It contains 3543 images, divided into 10 categories in total. We follow the official split.
+  • CheXpert [34] is a large dataset of chest X-rays and competition for automated chest x-ray interpretation, which features uncertainty labels and radiologist-labeled reference standard evaluation sets.
+It contains 224,316 chest radiographs if 65,240 patients, divided into 14 classes.
+We follow the official split.
+  • COVID-19-Radio [21, 66] contains 3616 COVID-19 positive cases along with 10,192 Normal, 6012 Lung Opacity (Non-COVID lung infection), and 1345 Viral Pneumonia 2D X-ray images and corresponding lung masks.
+We randomly split it to 2892 and 1220 for training and testing.
+  • IU-Xray [62] is a 2D chest X-ray datasets containing 7430 images in total. It is randomly split into 6674 for training and 756 for testing.
+  • LNDb [64] is developed as an external 3D dataset complimentary to LIDCIDRI.
+It contains 294 CT scans collected retrospectively at the Centro Hospitalar e Universitário de São João (CHUSJ) in Porto.
+We randomly split it into 235 for training and 59 for testing.
+  • PadChest [15] includes more than 160,000 2D chest X-ray images obtained from 67,000 patients that were interpreted and reported by radiologists.
+The reports were labeled with 174 different radiographic findings, 19 differential diagnoses and 104 anatomic locations organized as a hierarchical taxonomy and mapped onto standard Unified Medical Language System (UMLS) terminology.
+We use a subset of this dataset and randomly split it into 31242 for training and 7811 for testing.
+  • CC-CCII [81] contains 3777 3D volumns.
+All CT images are classified into novel coronavirus pneumonia (NCP) due to SARS-CoV-2 virus infection, common pneumonia and normal controls.
+We randomly split the dataset into 3021 for training and 756 for testing.
+  • RadChest [25] dataset is a large 3D medical imaging dataset developed by Duke University.
+The full dataset includes 35,747 chest CT scans from 19,661 adult patients.
+However, only 3,630 chest CT scans are available.
+There are 83 different labels in it.
+We randomly split it into 2904 for training and 726 for testing.
+  • Brain-Tumor [3] is a 2D dataset contains 7023 images of human brain MRI images which are classified into 4 classes: glioma, meningioma, no tumor and pituitary.
+We randomly split it into 5712 for training and 1311 for testing
+  • Brain-Tumor-17 [4] has 4449 real 2D MRI images of magnetic resonance imaging of the skull, in axial planes, weighted in T1, T1 with contrast and T2.
+It is classified into 6 classes: glioma, meningioma, neurocytoma, normal, other types of injuries and schwannoma.
+We randomly split the dataset into 3560 for training and 889 for testing.
+  • POCUS [12] is an chest ultrasound dataset containing more than 200 LUS videos labelled with a diagnostic outcome.
+It can be classified into 3 classes: COVID-19, pneumonia and regular.
+We treat this dataset as 2D images combination and randomly split into 2230 for training and 558 for testing.
+  • MURA [67] is a 2D X-ray dataset containing 14,863 musculoskeletal studies of the upper extremity, where each study contains one or more views and is manually labeled by radiologists as either normal or abnormal.
+We randomly split this dataset int 36806 for training and 3197 for testing.
+  • KneeMRI [70] is a 2D knee MRI dataset gathered retrospectively from exam records made on a Siemens Avanto 1.5T MR scanner.
+It contains 917 images, which can be classified into 3 classes.
+We randomly split this dataset into 733 for training and 184 for testing.
+  • CT-Kidney [35] is a 2D CT datasaet collected from PACS from different hospitals in Dhaka, Bangladesh where patients were already diagnosed with having a kidney tumor, cyst, normal or stone findings.
+It contains 12,446 unique data within it in which the cyst contains 3,709, normal 5,077, stone 1,377, and tumor 2,283.
+The train set and test set are randomly split as 9973/2473.
+4.3.3 Baselines For Comparison
+We show the comparison between our model and the existing SOTAs on the external public benchmark.
+In detail, under the fine-tuning setting, we compare with various specialized state-of-the-arts (SOTAs) for each datasets.
+Due to the large number of external datasets, we herein do not provide the details of the specific designs per method, while directly indicate the corresponding method by citation in the table, and more details of the designs can be found in the corresponding papers.
+In addition to the comparison with various specialized SOTAs, we also compare with other diagnosis foundation models in the zero-shot setting.
+In detail, we pick out related classes from our class set based on the target disease.
+It is possible that multiple similar classes in our dataset correspond to a single class in the external dataset due to the definition of class granularity, for example “Tumor” class in external dataset may be linked with “Brain Tumor”, “Chest Tumor” and so on in our class list.
+We treat the final prediction as positive if any class in the related class list is predicted as positive.
+Notably, as such union merging is carried out after thresholding the prediction scores, AUC score can not be calculated here.
+For comparison, we consider two types of baseline models:
+  • Classification Models.
+By training on a wide range of diagnosis data across various diseases, it enables to build a foundation model by simply optimizing the classification loss.
+Our work falls in this type, and another related work is RadIN [53].
+Similarly to our zero-shot adaptation pipeline, we also adopt RadIN to other datasets zero-shotly, by choosing a most related class from its original class list.
+  • Contrastive-learning Models.
+In addition to using classification labels, leveraging vision-language pairs, also enables to build up a foundation model through optimizing contrastive loss [18].
+Inspired by former works [72, 47], by treating the target disease class name as free-text, these models can also be adapted to zero-shot diagnosis leveraging the cross-modality similarity.
+We adopt the latest model, BiomedCLIP [82] form comparison.
+It is trained with large-scale image datasets sourced from PubMedCentral papers, and can handle various imaging modalities.
+However, it is worth emphasizing that both RadIN and BiomedCLIP can only handle 2D input.
+4.3.4 Metrics
+In this part, we will describe the evaluation metrics in detail.
+Note that, the following metrics can all be calculated per class.
+For multi-class cases, we all use macro-average on classes to report the scores by default.
+For example, “AUC” for multi-class classification denotes the “Macro-averaged AUC on classes”.
+We report all metrics in percent (%) in tables.
+AUC.
+Area Under Curve [14] denotes the area under ROC (receiver operating characteristic) curve.
+This has been widely used in medical diagnosis, due to its clinical meanings and robustness in unbalanced distribution.
+AP.
+Average Precision (AP) is calculated as the weighted mean of precisions at each threshold.
+Specifically, for each class, we rank all samples according to the prediction score, then shift the threshold to a series of precision-recall points and draw the precision-recall (PR) curve, AP equals the area under the curve.
+This score measures whether the unhealthy samples are ranked higher than healthy ones.
+We report Mean Average Precision (mAP), which is the average of AP of each class.
+F1, MCC and ACC.
+F1 score is the harmonic mean of the precision and recall.
+It is widely used in diagnosis tasks.
+ACC denotes the accuracy score.
+MCC [20] denotes the Matthews Correlation Coefficient metric.
+It ranges from $-1$ to $1$ and can be calculated as follow:
+$\text{MCC}=\frac{\text{TN}\times\text{TP}-\text{FN}\times\text{FP}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}$
+All the three metrics need a specific threshold to compute, and we choose one by maximizing F1 on the validation set following former papers [55, 77, 78, 83].
+Recall@FPR.
+We also report the recall scores at different false positive rate (FPR), i.e., sample points from class-wise ROC curves.
+Specifically, We report the Recall@0.01, Recall@0.05, Recall@0.1 score, denoting the recall scores at 0.01, 0.05, 0.1 FPR respectively.
+4.4 Implementation
+In this section, we will introduce implementation details such as the input and output during training and testing, model hyperparameters, and resource consumption information.
+As input to the model, the images are processed into shape of (H, W, D), where H and W represent the image height and width, both set to 256, and D represents the image depth, set to 32.
+To standardize the processing of both 2D and 3D images, we treat the 3D images as a stack of 2D slices.
+For images that do not conform to the specified dimensions, we employ bilinear interpolation for the height and width dimensions and uniform resampling for the depth dimension.
+Finally, we apply feature scaling for normalization.
+Based on the results of the ablation study for our model architecture, we select ResNet-50 and a transformer encoder for the visual encoder component.
+The features output by ResNet-50 have a shape of (2048, 8, 8).
+After applying average pooling, these features are passed through a transformer encoder to obtain features with a shape of (33, 2048).
+We choose the features of the first [cls] token for subsequent fusion among different scans.
+In the transformer encoder used for fusion, the number of heads is set to 4, the number of layers is set to 2, and the output feature dimension is 2048.
+Finally, through interaction with the knowledge embedding and a fully connected layer, we obtain the final prediction results with a shape of class numbers.
+For more detailed information, please refer to our Github repository.
+The training and testing for this work are conducted using Nvidia A100s (80G).
+During training, we apply image-level augmentation transformations.
+Additionally, considering that for each scan we have the manually picked key slice, we will also randomly replace the 3D scan with its 2D key slice.
+When both early fusion and data augmentation are activated, the GPU memory usage for training is approximately 46G with a batch size of 1.
+During inference, no data augmentation or random mix-up strategies are activated.
+Under this setting, GPU memory usage is approximately 7G with early fusion and a batch size of 1.
 5 CONCLUSION
-In this paper, we focus on solving the problem of multi-modal, multi-label long-tailed case-level diagnosis task.
-Specifically, we propose a new large-scale diagnosis dataset, namely, RP3D-DiagDS, with 39,026 cases (192,675 scans) labeled with detailed disorders covering 930 ICD-10-CM.
+In this work, we mainly focus on the diagnosis of general diseases (rare diseases included) based on radiology images, which can be formulated as a problem of multi-modal, multi-label long-tailed case-level diagnosis task.
+Specifically, we exploit the online data resources, constructing a new large-scale diagnosis dataset, namely, RP3D-DiagDS, with 40,936 cases (195,010 scans) labeled with detailed disorders covering 930 ICD-10-CM.
 On model design, we propose one unified architecture that supports both 2D and 3D input, together with a fusion module, to integrate information from multiple scans of one patient.
-Additionally, we adopt knowledge enhancement training strategy, leveraging the rich medical domain knowledge to improve the radiologic diagnosis performance.
-Our final train model also shows strong transferring ability to various external datasets regardless of their imaging modalities, shooting anatomies and target classes.
-We believe this work can serve as a transition phase between the specialist and generalist models, offering a more feasible and targeted playground for exploring sophisticated algorithms in academia labs, providing a unique opportunity for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs.
-10 Data and Code Availability
-Our dataset RP3D-DiagDS can be found in https://huggingface.co/datasets/QiaoyuZheng/RP3D-DiagDS and our codes can be found in https://github.com/qiaoyu-zheng/RP3D-Diag
+Additionally, we adopt knowledge enhancement training strategy, leveraging the rich medical domain knowledge to improve the radiological diagnosis performance.
+Our final model demonstrates strong performance on large-scale radiological disease diagnosis, and shows strong transferring ability to various external datasets regardless of their imaging modalities, shooting anatomies, and target classes.
+As a result, we show that publicly shared medical data on the Internet is a tremendous and valuable resource that can potentially support building a generalist model towards AI4healthcare.
+6 DATA AVAILABILITY
+Due to the Radiopaedia license  [^4] https://radiopaedia.org/licence?lang=us, we do not directly provide the original dataset.
+To obtain the data, readers must first get the permission from Radiopaedia’s founders, then we will share the download link.
+We also provide the model checkpoint in https://huggingface.co/QiaoyuZheng/RP3D-DiagModel following the apache-2.0 open-source license and all related result numbers can be reproduced based on the provided model checkpoint.
+7 CODE AVAILABILITY
+Our codes can be found in https://github.com/qiaoyu-zheng/RP3D-Diag under the MIT License, with all training, evaluation, fine-tuning, and zero-shot evaluation information.